### v1

# Imaginarity of quantum channels: Refinement and Alternative

Abstract
At present, quantum channels have been widely concerned, and many ways to quantify quantum channels have been proposed, which has led to the generation of many resource theories for quantum channels. We add strong monotonicity and convexity to the requirement of imaginarity measure of quantum channels to make the measure proper. We also introduce an alternative framework to simplify the process of verifying whether a quantifier is a proper measure. We present three imaginarity measures of quantum channels via on the robustness, the trace norm and entropy, respectively. Some properties are also given.
Keywords: quantum resource theories, quantum channels, imaginarity measures, alternative framework
1 Introduction
Imaginary numbers provide a broad perspective and additional methods for modern scientific research. Quantum resource theory can establish new connections among various research directions. With the development of quantum resource theory, the imaginarity resource theory [1] proposed in recent years has attracted increasing attention. Some examples of applications, the imaginarity plays a crucial role in state discrimination [2], the imaginarity operation is more advantageous to distinguish the quantum channel without the help of the auxiliary system [3]. Many imaginarity measures have been found for quantum states, the fidelity [2], the convex roof [4], the geometric [3] and some imaginarity measures induced by entropies [5, 6], some of which can establish relations due to their own mathematical properties [6, 7].
Just like quantum states, quantum channels also deserve attention. Some studies begin to investigate the properties of quantum channels [8, 9, 10, 11, 12]. Resource theories of quantum channels was established in [13], and many resource theories of quantum channels have been proposed, such as coherence of quantum channels [14], entanglement of quantum channels [15] and so on [16, 17]. Based on this, more measures for quantum channels have been found [18, 19, 20, 21].
In order to simplify the process of verifying whether a quantifier can be a coherence measure, alternative framework for quantifying coherence have been proposed [22] in the coherence resource theory. Similar alternative frameworks can be established in many other resource theories, such as imaginarity resource theory [7], coherence of quantum channels [23].
The concept of the imaginarity of quantum channels was first introduced in [24], and the definitions of free channels and free superchannels have been visited again recently [25]. It can be seen that the concept is worthing considing. Similar to the coherence of quantum channels, in this paper, we refine the imaginarity of quantum channels and introduce alternative framework to facilitate the verification of whether a quantifier can be a suitable measure. We also show three imaginarity measures of quantum channels. Some properties of the measure are discussed.
2 Refinement for the imaginarity of quantum channels
Let $H_{A}$ and $H_{B}$ be two Hilbert space with dimensions $|A|$ and $|B|$, and orthonormal basis $\left\{|j\rangle\right\}_{j=0}^{|A|-1}=\left\{|k\rangle\right\}_{k=0}^{|A|-1}$ and $\left\{|\alpha\rangle\right\}_{\alpha=0}^{|B|-1}=\left\{|\beta\rangle\right\}_{\beta=0}^{|B|-1}$. We assume $|a\rangle\langle a|=E_{a,a}$ without loss of generality, that is, it happens to be the a-th diagonal. Denote $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$ be the set of all density operators on $H_{A}$ and $H_{B}$, and $\mathcal{C}_{AB}$ be the set of all channels from $\mathcal{D}_{A}$ to $\mathcal{D}_{B}$. A quantum channel $\phi\in\mathcal{C}_{AB}$ can be represented by Choi matrix,
$J_{\phi}=\sum_{j,k}|j\rangle\langle k|\otimes\phi(|j\rangle\langle k|)=\sum_{j,k,\alpha,\beta}\phi_{j,k,\alpha,\beta}|j\rangle\langle k|\otimes|\alpha\rangle\langle\beta|$
with $\phi_{j,k,\alpha,\beta}=\langle\alpha|\phi(|j\rangle\langle k|)|\beta\rangle$.
The Choi state of $\phi$ is $\frac{J_{\phi}}{|A|}\triangleq\chi_{\phi}$. We call channel $\phi$ a real channel if $\chi_{\phi}$ is an real matrix, and denote $\mathcal{RC}_{AB}$ the set of all the real channels. The free channel is defined as the real channel. This is exactlt the free operation in imaginarity resource theory, thus it is appropriate to define the free channel by Choi state.
Let $\mathcal{SC}_{ABA^{\prime}B^{\prime}}$ be the sets of all superchannels from $\mathcal{C}_{AB}$ to $\mathcal{C}_{A^{\prime}B^{\prime}}$. It can be known that [14, 26] superchannel $\Theta$ also has the expression of Kraus operators $\left\{M_{m}\right\}_{m}$ with $\chi_{\Theta(\phi)}=\sum_{m}M_{m}\chi_{\phi}M_{m}^{\dagger}\triangleq\tilde{\Theta}(\chi_{\phi})$, where $\tilde{\Theta}$ is the channel corresponding to the superchannel. We will not distinguish the signs, $\Theta$ or $\tilde{\Theta}$, since it can be judged by the context.
We call a superchannel $\Theta:\mathcal{C}_{AB}\longrightarrow\mathcal{C}_{A^{\prime}B^{\prime}}$ a real superchannels if $\Theta(\phi)\in\mathcal{RC}_{A^{\prime}B^{\prime}}$ for any channels $\phi\in\mathcal{RC}_{AB}$. We denote the set of all the real superchannels by $\mathcal{RSC}_{ABA^{\prime}B^{\prime}}$. The free superchannel is defined as the real superchannel.
In [24], the author presented a framework for quantifying the imaginarity of quantum channels, they introduced the imaginarity measure $C(\phi)\triangleq C(\chi_{\phi})$ of quantum channels should satisfy the following conditions:
$\bm{(C_{1})}$. Faithfulness: $C(\phi)\geqslant 0$, and $C(\phi)=0$ if and only if $\phi\in\mathcal{RC}_{AB}$;
$\bm{(C_{2})}$. Monotonicity: $C(\phi)\geqslant C(\Theta(\phi))$ for any real superchannel $\Theta\in\mathcal{RSC}_{ABA^{\prime}B^{\prime}}$.
Strong monotonicity and convexity often have a great use and a proper quantum resource theory should include them. So we add these to the conditions above, and then the imaginarity measure $C$ of quantum channels becomes the more familiar and complete.
$\bm{(C_{3})}$. Strong monotonicity: Let $\Theta\in\mathcal{RSC}_{ABA^{\prime}B^{\prime}}$ with $\Theta(\cdot)=\sum_{m}M_{m}\cdot M_{m}^{\dagger}$, then $C(\phi)\geqslant\sum_{m}p_{m}C(\phi_{m})$ where $p_{m}=\text{Tr}(M_{m}\chi_{\phi}M_{m}^{\dagger})$ and $\chi_{\phi_{m}}=M_{m}\chi_{\phi}M_{m}^{\dagger}/p_{m}$
$\bm{(C_{4})}$. Convexity: For any $\left\{\phi_{m}\right\}\subset\mathcal{C}_{AB}$ and any probability distribution $\left\{p_{m}\right\}$, $C(\sum_{m}p_{m}\phi_{m})\leqslant\sum_{m}p_{m}C(\phi_{m})$.
We consider that a proper imaginarity measure of quantum channels should satisfy the condition $\bm{(C_{1})}$$\bm{(C_{4})}$. In this way, we can obtain some properties of the imaginarity measure $C$.
Property 1.
Monotonicity under composition with any real channels, that is,
$(a)$. Left composition: for any $\psi\in\mathcal{RC}_{BD}$ and any $\phi\in\mathcal{C}_{AB}$, we have: $C(\psi\circ\phi)\leqslant C(\phi)$.
$(b)$. Right composition: for any $\psi\in\mathcal{RC}_{AB}$ and any $\phi\in\mathcal{C}_{BD}$, we have: $C(\phi\circ\psi)\leqslant C(\phi)$.
Proof.
We only prove $(a)$, because $(b)$ is analogous.
For $\psi\in\mathcal{RC}_{AB}$ and $\phi\in\mathcal{C}_{BD}$, we consider a superchannel $\Theta_{\psi}$ with $\Theta_{\psi}=\psi\circ\phi$. Due to the condition $\bm{(C_{2})}$, we try to explain $\Theta\in\mathcal{RSC}_{ABAD}$. In fact, when superchannel $\Theta_{\psi}$ acts on any real channels, it is impossible to generate imaginarity resource since the channel $\psi$ is real. We can know that the superchannel $\Theta_{\psi}$ is a real superchannel from the definition. ∎
We can also see that if $\psi$ is invertible, $C(\psi\circ\phi)=C(\phi)$ holds.
Property 2.
The combination of imaginarity measures is still an imaginarity measure, that is, for any $\left\{C_{j}\right\}$ and any $\left\{q_{j}\right\}$ with $q_{j}\geqslant 0$, $C\triangleq\sum_{j}q_{j}C_{j}$ is an imaginarity measure .
Proof.
It is obvious that $C(\phi)\geqslant 0$ and due to $q_{j}\geqslant 0$, $C(\phi)=0$ is equivalent to every $C_{j}(\phi)=0$, certainly it indicate $\phi\in\mathcal{RC}$. For any real superchannel $\Theta$, we have $C(\phi)=\sum_{j}q_{j}C_{j}(\phi)\geqslant\sum_{j}q_{j}C_{j}(\Theta(\phi))=C(\Theta(\phi))$.
For $\Theta\in\mathcal{RSC}$ with $\Theta(\cdot)=\sum_{m}M_{m}\cdot M_{m}^{\dagger}$, the strong monotonicity holds:
$C(\phi)=\sum_{j}q_{j}C_{j}(\phi)\geqslant\sum_{j}q_{j}\sum_{m}p_{m}C_{j}(\phi_{m})=\sum_{m}p_{m}\sum_{j}q_{j}C_{j}(\phi_{m})=\sum_{m}p_{m}C(\phi_{m}).$
For any $\left\{\phi_{m}\right\}\subset\mathcal{C}_{AB}$ and any probability distribution $\left\{p_{m}\right\}$, the convexity holds:
$C(\sum_{m}p_{m}\phi_{m})=\sum_{j}q_{j}C_{j}(\sum_{m}p_{m}\phi_{m})\leqslant\sum_{j}q_{j}\sum_{m}p_{m}C_{j}(\phi_{m})=\sum_{m}p_{m}\sum_{j}q_{j}C_{j}(\phi_{m})=\sum_{m}p_{m}C(\phi_{m}).$
Therefore, $C$ is an imaginarity measure. ∎
Property 3.
If $C$ is an imaginarity measure for channel $\phi$, then, quantifier $\hat{C}$ as:
$\hat{C}(\phi)\triangleq\min_{\left\{q_{j},\phi_{j}\right\}}\sum_{j}q_{j}C(\phi_{j}),$
is a proper measure for $\phi$, where $\min_{\left\{q_{j},\phi_{j}\right\}}$ represents the min of all possible channels decompositions as $\phi=\sum_{j}q_{j}\phi_{j}$ with $q_{j}\geqslant 0$.
Proof.
It can be seen that $\hat{C}(\phi)\geqslant 0$. When $\hat{C}(\phi)=0$, we can take the best decomposition of $\phi$, expressed as $\phi=\sum_{k}q_{k}\phi_{k}$, so that $\hat{C}(\phi)=\sum_{k}q_{k}C(\phi_{k})$. We can get $C(\phi_{k})=0$ since $q\geqslant 0$, that is, $\phi_{k}\in\mathcal{RC}$, thus $\phi\in\mathcal{RC}$, the faithfulness holds.
We can notice that, for any $\left\{q_{j}\right\}$, let every $\phi_{j}=\phi$, we can get $\hat{C}(\phi)=\min_{\left\{q_{j},\phi_{j}\right\}}\sum_{j}q_{j}C(\phi_{j})\leqslant C(\phi)$. For any quantum channel $\phi$, we take the best decomposition of $\phi$, expressed as $\phi=\sum_{k}q_{k}\phi_{k}$, so that $\hat{C}(\phi)=\sum_{k}q_{k}C(\phi_{k})$. For $\Theta\in\mathcal{RSC}$ with $\Theta(\cdot)=\sum_{m}M_{m}\cdot M_{m}^{\dagger}$, we denote $a_{k,m}=\text{Tr}(M_{m}\chi_{\phi_{k}}M_{m}^{\dagger})$, $p_{m}=\text{Tr}(M_{m}\chi_{\phi}M_{m}^{\dagger})$, then
$\hat{C}(\phi)=\sum_{k}q_{k}C(\phi_{k})$$\geqslant$ $\sum_{k}q_{k}\sum_{m}a_{k,m}C(\frac{M_{m}\chi_{\phi_{k}}M_{m}^{\dagger}}{a_{k,m}})=\sum_{m}p_{m}\sum_{k}\frac{q_{k}a_{k,m}}{p_{m}}C(\frac{M_{m}\chi_{\phi_{k}}M_{m}^{\dagger}}{a_{k,m}})$
$\geqslant$ $\sum_{m}p_{m}\sum_{k}C(\frac{q_{k}a_{k,m}}{p_{m}}\frac{M_{m}\chi_{\phi_{k}}M_{m}^{\dagger}}{a_{k,m}})=\sum_{m}p_{m}C(\frac{\sum_{k}q_{k}M_{m}\chi_{\phi_{k}}M_{m}^{\dagger}}{p_{m}})$
$=$ $\sum_{m}p_{m}C(\frac{M_{m}\chi_{\phi}M_{m}^{\dagger}}{p_{m}})\geqslant\sum_{m}p_{m}\hat{C}(\frac{M_{m}\chi_{\phi}M_{m}^{\dagger}}{p_{m}}).$
The first inequality is due to the strong monotonicity of $C$, the second inequality is due to the convexity of $C$ and $\sum_{k}\frac{q_{k}a_{k,m}}{p_{m}}=\frac{\sum_{k}q_{k}\text{Tr}(M_{m}\chi_{\phi_{k}}M_{m}^{\dagger})}{\text{Tr}(M_{m}\chi_{\phi}M_{m}^{\dagger})}=1$.
To verify convexity of $\hat{C}$ for simplicity, we only consider two quantum channels $\phi_{1}$ and $\phi_{2}$, and define $\psi=a\phi_{1}+b\phi_{2}$ with $a+b=1$. We take the best decomposition of $\phi_{1}$ and $\phi_{2}$ respectively as $\phi_{1}=\sum_{m}q_{1,m}\phi_{1,m}$ and $\phi_{2}=\sum_{n}q_{2,n}\phi_{2,n}$ so that the equations hold:
$\begin{cases}\hat{C}(\phi_{1})=\sum_{m}q_{1,m}C(\phi_{1,m}),\\ \hat{C}(\phi_{2})=\sum_{n}q_{2,n}C(\phi_{2,n}).\end{cases}$
Then, $a\hat{C}(\phi_{1})+b\hat{C}(\phi_{2})=\sum_{m}aq_{1,m}C(\phi_{1,m})+\sum_{n}bq_{2,n}C(\phi_{2,n})$.
For ease of writing, if $m\neq n$, we supplement $\left\{0,\bm{0}\right\}$ for the less set so that both sets have the same quantity, that is, $a\hat{C}(\phi_{1})+b\hat{C}(\phi_{2})=\sum_{m}aq_{1,m}C(\phi_{1,m})+bq_{2,m}C(\phi_{2,m})$. Substitute the new symbol $\left\{q_{2m-1},\psi_{2m-1}\right\}$ and $\left\{q_{2m},\psi_{2m}\right\}$ for $\left\{aq_{1,m},\phi_{1,m}\right\}$ and $\left\{bq_{2,m},\phi_{2,m}\right\}$ respectively. We have
$a\hat{C}(\phi_{1})+b\hat{C}(\phi_{2})=\sum_{m}q_{2m-1}C(\psi_{2m-1})+q_{2m}C(\psi_{2m})\geqslant\min_{q_{j},\psi_{j}}\sum_{j}q_{j}C(\psi_{j})=\hat{C}(\psi)=\hat{C}(a\phi_{1}+b\phi_{2}).$
In this way, it can be known that $\hat{C}$ is a proper imaginarity measure. ∎
3 Alternative framework for the imaginarity of quantum channels
In the previous section we refined the framework. However, the condition $\bm{(C_{3})}$ and $\bm{(C_{4})}$ are difficult to verify in some cases. Here, inspired by the similar alternative framework in many resource theory [22, 23], we have the following result.
Theorem 1.
$\bm{(C_{3})}+\bm{(C_{4})}$ is equivalent to $\bm{(C_{5})}$ when condition $\bm{(C_{1})}$ and $\bm{(C_{2})}$ holds. The condition $\bm{(C_{5})}$ is:
$\bm{(C_{5})}$. Additivity: For any $p_{1}$, $p_{2}$ satisfied $p_{1}+p_{2}=1$, any $\phi_{1}\in\mathcal{C}_{AB_{1}}$ and $\phi_{2}\in\mathcal{C}_{AB_{2}}$, let $\phi=p_{1}\phi_{1}\oplus p_{2}\phi_{2}$, $\phi\in\mathcal{C}_{AB}$ with $|B|=|B_{1}|+|B_{2}|$, then $C(\phi)=p_{1}C(\phi_{1})+p_{2}C(\phi_{2})$ holds.
Proof.
The proof is long and we divide it into steps.
$\bm{(C_{3})}+\bm{(C_{4})}\Rightarrow\bm{(C_{5})}$:
(STEP 1).
Let $Q_{1}=|0\rangle\langle 0|+\cdots+||B_{1}|-1\rangle\langle|B_{1}|-1|+||B|\rangle\langle|B||+\cdots+||B|+|B_{1}|-1\rangle\langle|B|+|B_{1}|-1|+|(|A|-1)|B|\rangle\langle(|A|-1)|B||+\cdots+|(|A|-1)|B|+|B_{1}|-1\rangle\langle(|A|-1)|B|+|B_{1}|-1|$,
$Q_{2}=||B_{1}|\rangle\langle|B_{1}||+\cdots+||B|-1\rangle\langle|B|-1|+||B|+|B_{1}|\rangle\langle|B|+|B_{1}||+\cdots+||B|+|B|-1\rangle\langle|B|+|B|-1|+|(|A|-1)|B|+|B_{1}|\rangle\langle(|A|-1)|B|+|B_{1}||+\cdots+|(|A|-1)|B|+|B|-1\rangle\langle(|A|-1)|B|+|B|-1|$.
We can give an intuitive expression.
$Q_{1}=\overbrace{\begin{pmatrix}I_{|B_{1}|}&&&&\\ &O_{|B_{2}|}&&&\\ &&\ddots&&\\ &&&I_{|B_{1}|}&\\ &&&&O_{|B_{2}|}\end{pmatrix}}^{|A||B|}=\bigoplus\limits_{j=0}^{|A|-1}\begin{pmatrix}I_{|B_{1}|}&\\ &O_{|B_{2}|}\end{pmatrix},$
$Q_{2}=\underbrace{\begin{pmatrix}O_{|B_{1}|}&&&&\\ &I_{|B_{2}|}&&&\\ &&\ddots&&\\ &&&O_{|B_{1}|}&\\ &&&&I_{|B_{2}|}\end{pmatrix}}_{|A||B|}=\bigoplus\limits_{j=0}^{|A|-1}\begin{pmatrix}O_{|B_{1}|}&\\ &I_{|B_{2}|}\end{pmatrix},$
where $I_{a}$ and $O_{b}$ denote identity matrix and zero matrix with dimensions $a$ and $b$ respectively.
Due to $Q_{1}^{\dagger}Q_{1}+Q_{2}^{\dagger}Q_{2}=I_{|A||B|}$ and $Q_{1}Q_{2}=O_{|A||B|}$, let $\Theta(\cdot)=Q_{1}\cdot Q_{1}^{\dagger}+Q_{2}\cdot Q_{2}^{\dagger}$, we can know that $\Theta\in\mathcal{RSC}_{ABA^{\prime}B^{\prime}}$.
Under the requirement of $\bm{(C_{5})}$, the Choi state of $\phi$ is
$\chi_{\phi}=\sum_{j,k}|j\rangle\langle k|\otimes\phi(|j\rangle\langle k|)/|A|=\sum_{j,k}|j\rangle\langle k|\otimes\big{[}p_{1}\phi_{1}(|j\rangle\langle k|)\oplus p_{2}\phi_{2}(|j\rangle\langle k|\big{]}/|A|.$
Thus,
$Q_{1}\chi_{\phi}Q_{1}^{\dagger}=\sum_{j,k}|j\rangle\langle k|\otimes\big{[}p_{1}\phi_{1}(|j\rangle\langle k|)\oplus\bm{0}_{2}(|j\rangle\langle k|\big{]}/|A|.$
It’s obvious that $\text{Tr}(Q_{1}\chi_{\phi}Q_{1}^{\dagger})$ is exactly $p_{1}$. Let $\chi_{\tilde{\phi}_{1}}=Q_{1}\chi_{\phi}Q_{1}^{\dagger}/p_{1}=\sum_{j,k}|j\rangle\langle k|\otimes\big{[}\phi_{1}(|j\rangle\langle k|)\oplus\bm{0}_{2}(|j\rangle\langle k|\big{]}/|A|$. Similarly we obtain $\chi_{\tilde{\phi}_{2}}$. At this point, the requirement of $\bm{(C_{3})}$ has been satisfied, that is
$\begin{cases}\Theta\in\mathcal{RSC}_{ABA^{\prime}B^{\prime}}\text{ with }\Theta(\cdot)=Q_{1}\cdot Q_{1}^{\dagger}+Q_{2}\cdot Q_{2}^{\dagger},\\ p_{m}=\text{Tr}(Q_{m}\chi_{\phi}Q_{m}^{\dagger}),\chi_{\tilde{\phi}_{m}}=Q_{m}\chi_{\phi}Q_{m}^{\dagger}/p_{m},\end{cases}$
we have
$C(\chi_{\phi})\geqslant p_{1}C(\chi_{\tilde{\phi}_{1}})+p_{2}C(\chi_{\tilde{\phi}_{2}}).$ (1)
(STEP 2).
For any $\rho\in\mathcal{D}_{A}$, $\rho=\sum_{j,k}\rho_{j,k}|j\rangle\langle k|$, any $\phi\in\mathcal{C}_{AB}$, $\phi(\rho)$ can be represented by $J_{\phi}$, in order to do this, $\rho$ and $J_{\phi}$ need to be tensor index realigned:
$\begin{cases}J^{\prime}_{\phi}=\sum_{j,k}\langle jk|\otimes\phi(|j\rangle\langle k|),\\ \rho^{\prime}=\sum_{j,k}\rho_{j,k}|jk\rangle,\end{cases}$
then $\phi(\rho)=J^{\prime}_{\phi}\rho^{\prime}=\sum_{j,k}\rho_{j,k}\phi(|j\rangle\langle k|)$. Under the requirement of $(C_{5})$, we have
$\phi(\rho)$ $=$ $\sum_{j,k}\rho_{j,k}\phi(|j\rangle\langle k|)=\sum_{j,k}\rho_{j,k}\big{[}p_{1}\phi_{1}(|j\rangle\langle k|)\oplus p_{2}\phi_{2}(|j\rangle\langle k|)\big{]}$
$=$ $p_{1}\sum_{j,k}\rho_{j,k}\big{[}\phi_{1}(|j\rangle\langle k|)\oplus\bm{0}_{2}(|j\rangle\langle k|)\big{]}+p_{2}\sum_{j,k}\rho_{j,k}\big{[}\bm{0}_{1}(|j\rangle\langle k|)\oplus\phi_{2}(|j\rangle\langle k|)\big{]}$
$=$ $p_{1}\tilde{\phi}_{1}(\rho)+p_{2}\tilde{\phi}_{2}(\rho),$
that is $\phi=p_{1}\tilde{\phi}_{1}+p_{2}\tilde{\phi}_{2}$. From $\bm{(C_{4})}$ We can know that
$C(\phi)\leqslant p_{1}C(\tilde{\phi}_{1})+p_{2}C(\tilde{\phi}_{2}).$ (2)
(STEP 3).
From (1) in (STEP 1) and (2) in (STEP 2), we obtain $C(\phi)=p_{1}C(\tilde{\phi}_{1})+p_{2}C(\tilde{\phi}_{2})$. Next we need to prove $C(\phi_{1})=C(\tilde{\phi}_{1})$, that is, we need to prove $C(\chi_{\phi_{1}})=C(\chi_{\tilde{\phi}_{1}})$.
First, we specify the range of index $i$ and $j$, let $i=0,1,\cdots,|B|-1$, $j=0,1,\cdots,|B_{1}|-1$. Let $K_{0}^{1}$ satisfies $\langle i|K_{0}^{1}|j\rangle=\delta_{i,j}$. An intuitive expression of $K_{0}^{1}$ is
$K_{0}^{1}=\left.\begin{pmatrix}I_{|B_{1}|}\\ \hat{O}\end{pmatrix}\right\}|B|,$
where $\hat{O}$ denotes zero matrix with dimension $|B_{2}|\times|B_{1}|$.
Let $K_{n}^{2}$ satisfy $\langle j|K_{n}^{2}|i\rangle=\delta_{i,j+n|B_{1}|}$, $n=0,1,\cdots,\left\lceil|B_{2}|/|B_{1}|\right\rceil$. $\left\lceil a\right\rceil$ is the ceil of $a$. We can give an example to make this easier to understand. See Example 1 after this proof.
Let
$\mathcal{K}_{0}^{1}=\bigoplus_{j=0}^{|A|-1}K_{0}^{1}=\begin{pmatrix}K_{0}^{1}&&\\ &\ddots&\\ &&K_{0}^{1}\end{pmatrix},\mathcal{K}_{n}^{2}=\bigoplus_{j=0}^{|A|-1}K_{n}^{2}=\begin{pmatrix}K_{n}^{2}&&\\ &\ddots&\\ &&K_{n}^{2}\end{pmatrix},$
we can verify $(\mathcal{K}_{0}^{1})^{\dagger}\mathcal{K}_{0}^{1}=I_{|B_{1}|}$ and $\sum_{j=0}^{\left\lceil|B_{2}|/|B_{1}|\right\rceil}(\mathcal{K}_{n}^{2})^{\dagger}\mathcal{K}_{n}^{2}=I_{|B|}$. Thus $\Theta_{1}\in\mathcal{RSC}_{AB_{1}AB}$ with $\Theta_{1}(\cdot)=\mathcal{K}_{0}^{1}\cdot(\mathcal{K}_{0}^{1})^{\dagger}$, $\Theta_{2}\in\mathcal{RSC}_{ABAB_{1}}$ with $\Theta_{2}(\cdot)=\sum_{j=0}^{\left\lceil|B_{2}|/|B_{1}|\right\rceil}\mathcal{K}_{n}^{2}\cdot(\mathcal{K}_{n}^{2})^{\dagger}$. We can also obtain $\Theta_{1}(\chi_{\phi_{1}})=\chi_{\tilde{\phi}_{1}}$ and $\Theta_{2}(\chi_{\tilde{\phi}_{1}})=\chi_{\phi_{1}}$.
Combining with $\bm{(C_{2})}$, the above implies that the following inequality holds:
$C(\chi_{\phi_{1}})=C(\Theta_{2}(\chi_{\tilde{\phi}_{1}}))\leqslant C(\chi_{\tilde{\phi}_{1}})=C(\Theta_{1}(\chi_{\phi_{1}}))\leqslant C(\chi_{\phi_{1}}).$
As a result, $C(\chi_{\phi_{1}})=C(\chi_{\tilde{\phi}_{1}})$. It follows from a similar discussion that $C(\chi_{\phi_{2}})=C(\chi_{\tilde{\phi}_{2}})$.
Up to now, we prove $\bm{(C_{5})}$ holds: $C(\phi)=p_{1}C(\phi_{1})+p_{2}C(\phi_{2})$.
$\bm{(C_{5})}\Rightarrow\bm{(C_{3})}+\bm{(C_{4})}$:
(STEP 4).
For $\Theta\in\mathcal{RSC}_{ABA^{\prime}B^{\prime}}$ with $\Theta(\cdot)=\sum_{m}M_{m}\cdot M_{m}^{\dagger}$, we introduce an auxiliary system $S$ with dimension $N$ and orthonormal basis $\left\{|n\rangle\right\}_{n=0}^{N-1}$. The system $AB$ formed by Choi state and the auxiliary system $S$ form a combined system $ABS$, of which the basis is $\left\{|j\alpha\rangle\otimes|n\rangle\right\}$ and the whole system is initially in the state $\chi_{\phi}\otimes|0\rangle\langle 0|\triangleq\chi_{\phi}^{ABS}$.
Let $U_{n}=\sum_{k=0}^{N-1}|(k+n)\text{ mod }N\rangle\langle k|$, then $\Theta^{ABS}(\cdot)\in\mathcal{RSC}_{ABSA^{\prime}B^{\prime}S}$ with $\Theta^{ABS}(\cdot)=\sum_{n}(M_{n}\otimes U_{n})\cdot(M_{n}\otimes U_{n})^{\dagger}$. Thus
$\Theta^{ABS}(\chi_{\phi}^{ABS})=\sum_{n}(M_{n}\otimes U_{n})(\chi_{\phi}\otimes|0\rangle\langle 0|)(M_{n}\otimes U_{n})^{\dagger}=\sum_{n}M_{n}\chi_{\phi}M_{n}^{\dagger}\otimes|n\rangle\langle n|\triangleq\sum_{n}p_{n}\chi_{\phi_{n}}\otimes|n\rangle\langle n|,$
where $p_{n}=\text{Tr}(M_{n}\chi_{\phi}M_{n}^{\dagger})$ and $\chi_{\phi_{n}}=M_{n}\chi_{\phi}M_{n}^{\dagger}/p_{n}$.
We can know that $\sum_{n}p_{n}=1$, $\phi_{n}\in\mathcal{C}_{AB_{n}}$ with $|B|=\sum_{n}|B_{n}|$. From $\bm{(C_{5})}$, we get
$\begin{cases}C(\chi_{\phi}^{ABS})=C(\chi_{\phi}\otimes|0\rangle\langle 0|)=C(\chi_{\phi}),\\ C(\Theta^{ABS}(\chi_{\phi}^{ABS}))=C(\sum_{n}p_{n}\chi_{\phi_{n}}\otimes|n\rangle\langle n|)=C(\bigoplus_{n}p_{n}\chi_{\phi_{n}})=\sum_{n}p_{n}C(\chi_{\phi_{n}}).\end{cases}$
From $\bm{(C_{2})}$, we can obtain
$C(\chi_{\phi})=C(\chi_{\phi}^{ABS})\geqslant C(\Theta^{ABS}(\chi_{\phi}^{ABS}))=\sum_{n}p_{n}C(\chi_{\phi_{n}}),$
that is, condition $\bm{(C_{3})}$ holds.
(STEP 5).
For any $\left\{\phi_{m}\right\}\subset\mathcal{C}_{AB}$ and any probability distribution $\left\{p_{m}\right\}$, we construct the similar auxiliary system $S$ with orthonormal basis $\left\{|m\rangle\right\}_{m}$ and define $\chi_{\phi}^{ABS}=\sum_{m}p_{m}\chi_{\phi_{m}}\otimes|m\rangle\langle m|$. Due to $\sum_{m}(I_{|A||B|}\otimes|0\rangle\langle m|)^{\dagger}(I_{|A||B|}\otimes|0\rangle\langle m|)=I_{|A||B||S|}$, $\Theta^{ABS}\in\mathcal{RSC}_{ABSA^{\prime}B^{\prime}S}$ with $\Theta^{ABS}(\cdot)=\sum_{m}(I_{|A||B|}\otimes|0\rangle\langle m|)\cdot(I_{|A||B|}\otimes|0\rangle\langle m|)^{\dagger}$. Then
$\Theta^{ABS}(\chi_{\phi}^{ABS})=\sum_{m}(I_{|A||B|}\otimes|0\rangle\langle m|)\chi_{\phi}^{ABS}(I_{|A||B|}\otimes|0\rangle\langle m|)^{\dagger}=\sum_{m}p_{m}\chi_{\phi_{m}}\otimes|0\rangle\langle 0|.$
We know from $\bm{(C_{2})}$ and $\bm{(C_{5})}$ that the following inequality holds:
$\sum_{m}p_{m}C(\chi_{\phi_{m}})=C(\chi_{\phi}^{ABS})\geqslant C(\Theta^{ABS}(\chi_{\phi}^{ABS}))=C(\sum_{m}p_{m}\chi_{\phi_{m}}),$
that is, condition $\bm{(C_{4})}$ holds.
Therefore, the proof is finished. ∎
Example 1.
Suppose $|B_{1}|=2$ and $|B_{2}|=3$, then $i=0,1,\cdots,4$, $j=0,1$, $n=0,1,2$ and
$K_{0}^{2}=\begin{pmatrix}1&0&0&0&0\\ 0&1&0&0&0\end{pmatrix}\triangleq\begin{pmatrix}I_{|B_{1}|}&O_{|B_{1}|}&\tilde{O}\end{pmatrix},$
$K_{1}^{2}=\begin{pmatrix}0&0&1&0&0\\ 0&0&0&1&0\end{pmatrix}\triangleq\begin{pmatrix}O_{|B_{1}|}&I_{|B_{1}|}&\tilde{O}\end{pmatrix},$
$K_{2}^{2}=\begin{pmatrix}0&0&0&0&1\\ 0&0&0&0&0\end{pmatrix}\triangleq\begin{pmatrix}O_{|B_{1}|}&O_{|B_{1}|}&\tilde{I}\end{pmatrix},$
where $\tilde{O}$ denotes the first $(|B_{2}|\text{ mod }|B_{1}|)$ columns of $O_{|B_{1}|}$, same does $\tilde{I}$.
It is often difficult to verify whether a measure has strong monotonicity and convexity directly, which may be easier after using Theorem 1. When verifying whether a quantifier is a proper measure, we can choose verifying $\bm{(C_{1})}$$\bm{(C_{4})}$ or verifying $\bm{(C_{1})}+\bm{(C_{2})}+\bm{(C_{5})}$.
Under the requirement of $\bm{(C_{5})}$, that is, for any $p_{1}$, $p_{2}$ satisfied $p_{1}+p_{2}=1$, any $\phi_{1}\in\mathcal{C}_{AB_{1}}$ and $\phi_{2}\in\mathcal{C}_{AB_{2}}$, let $\phi=p_{1}\phi_{1}\oplus p_{2}\phi_{2}$, $\phi\in\mathcal{C}_{AB}$ with $|B|=|B_{1}|+|B_{2}|$. We have $\chi_{\phi_{1}}$ and $\chi_{\phi_{2}}$, and we define $\chi_{\tilde{\phi}_{1}}$ and $\chi_{\tilde{\phi}_{2}}$ below.
$\begin{cases}\chi_{\phi_{1}}=\sum_{j,k}|j\rangle\langle k|\otimes\phi_{1}(|j\rangle\langle k|)/|A|,\\ \chi_{\phi_{2}}=\sum_{j,k}|j\rangle\langle k|\otimes\phi_{2}(|j\rangle\langle k|)/|A|,\\ \chi_{\tilde{\phi}_{1}}=\sum_{j,k}|j\rangle\langle k|\otimes\big{[}\phi_{1}(|j\rangle\langle k|)\oplus\bm{0}_{2}(|j\rangle\langle k|\big{]}/|A|,\\ \chi_{\tilde{\phi}_{2}}=\sum_{j,k}|j\rangle\langle k|\otimes\big{[}\bm{0}_{1}(|j\rangle\langle k|)\oplus\phi_{2}(|j\rangle\langle k|\big{]}/|A|.\end{cases}$
Thus the following property holds.
Property 4.
$C$ is a proper imaginarity measure for quantum channels, the equation holds:
$C(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})=C(p_{1}\chi_{\tilde{\phi}_{1}}+p_{2}\chi_{\tilde{\phi}_{2}})=C(\chi_{\phi}).$ (3)
The proof of Property 4 is obvious, here is a helpful example.
Example 2.
The following matrix transformations hold and they are invertible.
$\begin{pmatrix}a_{11}&a_{12}&&&&\\ a_{21}&a_{22}&&&&\\ &&a_{33}&&&\\ &&&b_{11}&b_{12}&\\ &&&b_{21}&b_{22}&\\ &&&&&b_{33}\end{pmatrix}\xrightarrow{K_{1}(\cdot)K_{1}^{\dagger}}\begin{pmatrix}a_{11}&&a_{12}&&&\\ &a_{33}&&&&\\ a_{21}&&a_{22}&&&\\ &&&b_{22}&b_{21}&\\ &&&b_{12}&b_{11}&\\ &&&&&b_{33}\end{pmatrix}\xrightarrow{K_{2}(\cdot)K_{2}^{\dagger}}\begin{pmatrix}a_{11}&&a_{12}&&&\\ &b_{11}&&b_{12}&&\\ a_{21}&&a_{22}&&&\\ &b_{21}&&b_{22}&&\\ &&&&a_{33}&\\ &&&&&b_{33}\\ \end{pmatrix},$
where
$K_{1}=\begin{pmatrix}1&&&&&\\ &&1&&&\\ &1&&&&\\ &&&&1&\\ &&&1&&\\ &&&&&1\\ \end{pmatrix},K_{2}=\begin{pmatrix}1&&&&&\\ &&&&1&\\ &&1&&&\\ &&&1&&\\ &1&&&&\\ &&&&&1\\ \end{pmatrix}.$
Due to $K^{\dagger}_{j}K_{j}=I$, then the invertible $\Theta_{j}\in\mathcal{RSC}$ with $\Theta_{j}(\cdot)=K_{j}\cdot K_{j}^{\dagger}$, $j=1,2$.
Just like $\Theta_{1}$ and $\Theta_{2}$ in this example, we can construct multiple invertible channels such that
$\Theta_{1}(\Theta_{2}(\cdots\Theta_{k}(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})\cdots))=p_{1}\chi_{\tilde{\phi}_{1}}+p_{2}\chi_{\tilde{\phi}_{2}}.$
In this way, we can know that the equation (3) holds from the condition $\bm{(C_{2})}$.
4 Imaginarity measures of quantum channels
The Choi theorem [27] tell us that the Choi matrix of the channel $\phi$ is positive semidefinite, so that $\chi_{\phi}\geqslant 0$, we can see that, for any quantum channel $\phi$ with $\phi(\cdot)=\sum_{n}K_{n}\cdot K_{n}^{\dagger}$, its Choi state $\chi_{\phi}=\sum_{j,k}|j\rangle\langle k|\otimes\phi(|j\rangle\langle k|)/|A|$, thus the trace:
$\text{Tr}(\chi_{\phi})=\text{Tr}\Big{[}\sum_{j,k}|j\rangle\langle k|\otimes\phi(|j\rangle\langle k|)/|A|\Big{]}=\frac{1}{|A|}\sum_{j}\text{Tr}\Big{[}\phi(|j\rangle\langle j|)\Big{]}=\frac{1}{|A|}\sum_{j}1=1.$
In this way, we can embed Choi states into quantum states:
$\mathcal{C}_{AB}\hookrightarrow\mathcal{D}_{AB}.$
It seems then that many imaginarity measures for quantum states can be equally applied to quantum channels. Thus in this section, we introduce three imaginarity measures of quantum channels, that is, measures via on the robustness, the trace norm and the sandwiched Rényi relative entropy, respectively. We can exactly verify these three measures in two ways.
4.1 Measure via on the robustness
The Measure via on the robustness is appeared in many resource theories [1, 18]. Similar to these, we define the imaginarity measure of quantum channels as follow.
$C_{r}(\phi)=\min_{\tilde{\phi}\in\mathcal{RC}_{AB}}\left\{s\geqslant 0:\frac{\chi_{\phi}+s\chi_{\tilde{\phi}}}{1+s}\triangleq\chi_{\psi},\psi\in\mathcal{RC}_{AB}\right\}$
We have the following result.
Theorem 2.
$C_{r}(\phi)$ is a proper imaginarity measure of quantum channels.
Before our proof, we need to introduce pseudomixture [28, 29, 30, 7] of $\chi_{\phi}$, that is,
$\chi_{\phi}=[1+C_{r}(\chi_{\phi})]\chi_{\phi_{1}}-C_{r}(\chi_{\phi})\chi_{\phi_{2}},$
where $\phi_{1}\in\mathcal{RC}_{AB}$, and $\phi_{2}$ is the optimal channel of the $\min$ in the definition of $C_{r}(\phi)$.
Proof.
We only verify $\bm{(C_{1})}$+$\bm{(C_{3})}$+$\bm{(C_{4})}$ since $\bm{(C_{2})}$ can be derived.
$\bm{(C_{1})}$:
Obviously $C_{r}(\phi)\geqslant 0$. If $C_{r}(\phi)=0$, that means $\frac{\chi_{\phi}+0}{1+0}=\chi_{\psi}$, that is $\phi=\psi\in\mathcal{RC}_{AB}$. The converse is certainly true.
$\bm{(C_{3})}$:
Let $\Theta\in\mathcal{RSC}_{ABA^{\prime}B^{\prime}}$ with $\Theta(\cdot)=\sum_{m}M_{m}\cdot M_{m}^{\dagger}$, for any quantum channel $\phi$ with its pseudomixture, under the requirement of $\bm{(C_{3})}$, $p_{m}=\text{Tr}(M_{m}\chi_{\phi}M_{m}^{\dagger})$ and $\chi_{\phi_{m}}=M_{m}\chi_{\phi}M_{m}^{\dagger}/p_{m}$. Then we obtain
$M_{m}\chi_{\phi}M_{m}^{\dagger}=[1+C_{r}(\chi_{\phi})]M_{m}\chi_{\phi_{1}}M_{m}^{\dagger}-C_{r}(\chi_{\phi})M_{m}\chi_{\phi_{2}}M_{m}^{\dagger}.$
Let
$\begin{cases}\chi_{\phi_{1}}^{m}=\frac{1}{1+s_{m}}\frac{1}{p_{m}}[1+C_{r}(\chi_{\phi})]M_{m}\chi_{\phi_{1}}M_{m}^{\dagger},\\ \chi_{\phi_{2}}^{m}=\frac{1}{s_{m}}\frac{1}{p_{m}}C_{r}(\chi_{\phi})M_{m}\chi_{\phi_{2}}M_{m}^{\dagger},\end{cases}$
where $s_{m}=\frac{1}{p_{m}}C(\chi_{\phi})\text{Tr}(M_{m}\chi_{\phi_{1}}M_{m}^{\dagger})$.
We get $\chi_{\phi_{m}}=(1+s_{m})\chi_{\phi_{1}}^{m}-s_{m}\chi_{\phi_{2}}^{m}$, that is, $\frac{\chi_{\phi_{m}}+s_{m}\chi_{\phi_{2}}^{m}}{1+s_{m}}=\chi_{\phi_{1}}^{m}$. Thus we get $C_{r}(\chi_{\phi_{m}})\leqslant s_{m}$ from definition, that is, $p_{m}C_{r}(\chi_{\phi_{m}})\leqslant C(\chi_{\phi})\text{Tr}(M_{m}\chi_{\phi_{1}}M_{m}^{\dagger})$.
Therefore, we obtain
$\sum_{m}p_{m}C_{r}(\chi_{\phi_{m}})$$\leqslant$ $\sum_{m}C(\chi_{\phi})\text{Tr}(M_{m}\chi_{\phi_{1}}M_{m}^{\dagger})$
$=$ $C(\chi_{\phi})\sum_{m}\text{Tr}(M_{m}^{\dagger}M_{m}\chi_{\phi_{1}})$
$=$ $C(\chi_{\phi})\text{Tr}(\sum_{m}M_{m}^{\dagger}M_{m}\chi_{\phi_{1}})$
$=$ $C(\chi_{\phi})$
This means the condition $\bm{(C_{3})}$ holds.
$\bm{(C_{4})}$:
For simplicity, we only consider two quantum channels $\phi_{1}$ and $\phi_{2}$, and define $\phi=p\phi_{1}+(1-p)\phi_{2}$ with any $p\in(0,1)$, that is, $\chi_{\phi}=p\chi_{\phi_{1}}+(1-p)\chi_{\phi_{2}}$. The measurements $C_{r}$ for $\phi_{1}$ and $\phi_{2}$ are $C_{r}(\chi_{\phi_{1}})$ and $C_{r}(\chi_{\phi_{2}})$. Let $s_{p}=pC_{r}(\chi_{\phi_{1}})+(1-p)C_{r}(\chi_{\phi_{2}})$.
Consider the pseudomixtures of $\chi_{\phi_{1}}$ and $\chi_{\phi_{2}}$:
$\chi_{\phi_{1}}=[1+C_{r}(\chi_{\phi_{1}})]\chi_{\phi_{1,1}}-C_{r}(\chi_{\phi_{1}})\chi_{\phi_{1,2}}\mbox{ and }\chi_{\phi_{2}}=[1+C_{r}(\chi_{\phi_{2}})]\chi_{\phi_{2,1}}-C_{r}(\chi_{\phi_{2}})\chi_{\phi_{2,2}},$
then let
$\begin{cases}\chi_{\psi}=\frac{p[1+C_{r}(\chi_{\phi_{1}})]\chi_{\phi_{1,1}}+(1-p)[1+C_{r}(\chi_{\phi_{2}})]\chi_{\phi_{2,1}}}{1+s_{p}},\\ \chi_{\tilde{\phi}}=\frac{pC_{r}(\chi_{\phi_{1}})\chi_{\phi_{1,2}}+(1-p)C_{r}(\chi_{\phi_{2}})\chi_{\phi_{2,2}}}{s_{p}}.\end{cases}$
We have $\chi_{\phi}=(1+s_{p})\chi_{\psi}-s_{p}\chi_{\tilde{\phi}}$, that is, $\chi_{\psi}=\frac{\chi_{\phi}+s_{p}\chi_{\tilde{\phi}}}{1+s_{p}}$, thus we get $s_{p}\geqslant C_{r}(\chi_{\phi})$.
Therefore, $pC_{r}(\chi_{\phi_{1}})+(1-p)C_{r}(\chi_{\phi_{2}})=s_{p}\geqslant C_{r}(\chi_{\phi})=C(p\chi_{\phi_{1}}+(1-p)\chi_{\phi_{2}})$. The condition $\bm{(C_{4})}$ holds. ∎
4.2 Measure via on the trace norm
To obtain the channel is real, we can consider the trace norm:
$C_{t}(\phi)=||\chi_{\phi}-\chi_{\phi}^{*}||_{\text{Tr}}=\text{Tr}|\chi_{\phi}-\chi_{\phi}^{*}|.$
where $\chi_{\phi}^{*}$ means the conjugate of $\chi_{\phi}$ and $|\chi_{\phi}-\chi_{\phi}^{*}|=\sqrt{(\chi_{\phi}-\chi_{\phi}^{*})^{\dagger}(\chi_{\phi}-\chi_{\phi}^{*})}$.
Theorem 3.
$C_{t}(\phi)$ is a proper imaginarity measure of quantum channels.
Proof.
The condition $\bm{(C_{1})}$ holds due to the definition of the trace norm.
There is $||\chi_{\phi}-\chi_{\phi}^{*}||_{\text{Tr}}\geqslant||\Theta(\chi_{\phi})-\Theta(\chi_{\phi}^{*})||_{\text{Tr}}$ for any CPTP map [31], and it certainly holds for real superchannels. Since $\Theta$ is a real superchannel, we can know that $[\Theta(\chi_{\phi})]^{*}=\Theta(\chi_{\phi}^{*})$ due to
$[\Theta(\chi_{\phi})]^{*}=\left[\sum_{m}M_{m}\chi_{\phi}M_{n}^{\dagger}\right]^{*}=\sum_{m}M_{m}\chi_{\phi}^{*}M_{n}^{\dagger}=\Theta(\chi_{\phi}^{*}).$
Thus $\bm{(C_{2})}$ is satisfied.
Under the requirement of $\bm{(C_{5})}$, from Property 4, we have:
$C_{t}(\chi_{\phi})$ $=$ $C_{t}(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})=||(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})-(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})^{*}||_{\text{Tr}}$
$=$ $||(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})-(p_{1}\chi_{\phi_{1}}^{*}\oplus p_{2}\chi_{\phi_{2}}^{*})||_{\text{Tr}}=||p_{1}(\chi_{\phi_{1}}-\chi_{\phi_{1}}^{*})\oplus p_{2}(\chi_{\phi_{2}}-\chi_{\phi_{2}}^{*})||_{\text{Tr}}$
$=$ $p_{1}||(\chi_{\phi_{1}}-\chi_{\phi_{1}}^{*})||_{\text{Tr}}+p_{2}||(\chi_{\phi_{2}}-\chi_{\phi_{2}}^{*})||_{\text{Tr}}=p_{1}C_{t}(\chi_{\phi_{1}})+p_{2}C_{t}(\chi_{\phi_{2}}).$
Therefore, $C_{r}(\phi)$ satisfies $\bm{(C_{1})}$+$\bm{(C_{2})}$+$\bm{(C_{5})}$. ∎
4.3 Measure via on the sandwiched Rényi relative entropy
The sandwiched Rényi relative entropy [32] defined for quantum states with the parameter $\alpha$ as follow:
$D_{\alpha}(\rho||\sigma)=\frac{1}{\alpha-1}\text{log }\text{Tr}(\sigma^{\frac{1-\alpha}{2\alpha}}\rho\sigma^{\frac{1-\alpha}{2\alpha}})^{\alpha},$
where Now, we will introduce the imaginarity measure of quantum channels via on the sandwiched Rényi relative entropy, that is,
$C_{\alpha}(\phi)=1-\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha},$
where $\frac{1}{2}\leqslant\alpha<1$. We have the following results.
Theorem 4.
$C_{\alpha}(\phi)$ is a proper imaginarity measure of quantum channels.
Proof.
It is shown [32] that $D_{\alpha}(\rho||\sigma)\geqslant 0$ when $\frac{1}{2}\leqslant\alpha<1$ and equation holds if and only if $\rho=\sigma$, that is, $\text{Tr}(\sigma^{\frac{1-\alpha}{2\alpha}}\rho\sigma^{\frac{1-\alpha}{2\alpha}})^{\alpha}\leqslant 1$. Thus we consider Choi state as a quantum state and then we have, $C_{\alpha}(\phi)\geqslant 0$. The equation holds if and only if $\chi_{\phi}=\chi_{\phi}^{*}$. The condition $\bm{(C_{1})}$ is satisfied.
From [32], we have:
$D_{\alpha}(\chi_{\phi}||\chi_{\phi}^{*})\geqslant D_{\alpha}\left[\Theta(\chi_{\phi})||\Theta(\chi_{\phi}^{*})\right]$
$\iff$ log $\text{log }\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}\leqslant\text{log }\text{Tr}\left\{[\Theta(\chi_{\phi}^{*})]^{\frac{1-\alpha}{2\alpha}}\Theta(\chi_{\phi})[\Theta(\chi_{\phi}^{*})]^{\frac{1-\alpha}{2\alpha}}\right\}^{\alpha}$
$\iff$ $\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}\leqslant\text{Tr}\left\{[\Theta(\chi_{\phi}^{*})]^{\frac{1-\alpha}{2\alpha}}\Theta(\chi_{\phi})[\Theta(\chi_{\phi}^{*})]^{\frac{1-\alpha}{2\alpha}}\right\}^{\alpha}$
$\iff$ $\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}\leqslant\text{Tr}\left\{([\Theta(\chi_{\phi})]^{*})^{\frac{1-\alpha}{2\alpha}}\Theta(\chi_{\phi})([\Theta(\chi_{\phi})]^{*})^{\frac{1-\alpha}{2\alpha}}\right\}^{\alpha}$
$\iff$ $C_{\alpha}(\phi)\geqslant C_{\alpha}\left[\Theta(\phi)\right].$
Thus $\bm{(C_{2})}$ holds.
Under the requirement of $\bm{(C_{5})}$, we have:
$\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}$ $=$ $\text{Tr}\left\{\left[(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})^{*}\right]^{\frac{1-\alpha}{2\alpha}}(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})\left[(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})^{*}\right]^{\frac{1-\alpha}{2\alpha}}\right\}^{\alpha}$
$=$ $\text{Tr}\left\{\bigoplus_{j=1}^{2}p_{j}^{\frac{1-\alpha}{2\alpha}}p_{j}p_{j}^{\frac{1-\alpha}{2\alpha}}(\chi_{\phi_{j}}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi_{j}}(\chi_{\phi_{j}}^{*})^{\frac{1-\alpha}{2\alpha}}\right\}^{\alpha}$
$=$ $\sum_{j=1}^{2}p_{j}\text{Tr}\left[(\chi_{\phi_{j}}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi_{j}}(\chi_{\phi_{j}}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}$
In this way, we obtain:
$C_{\alpha}(\phi)=1-\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}=\sum_{j=1}^{2}p_{j}\left\{1-\text{Tr}\left[(\chi_{\phi_{j}}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi_{j}}(\chi_{\phi_{j}}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}\right\}=p_{1}C_{\alpha}(\phi_{1})+p_{2}C_{\alpha}(\phi_{2})$
The condition $\bm{(C_{5})}$ holds. ∎
5 Conclusion
We refine the framework of imaginarity of quantum channels which originally only has $\bm{(C_{1})}$ and $\bm{(C_{2})}$ for its measure. We add two conditions, strong monotonicity $\bm{(C_{3})}$ and convexity $\bm{(C_{4})}$. Besides, we give some properties of imaginarity measure of quantum channels. Sometimes it is not easy to verify strong monotonicity and convexity, so we introduce an alternative framework, that is, $\bm{(C_{3})}+\bm{(C_{4})}$ is equivalent to $\bm{(C_{5})}$ when $\bm{(C_{1})}$ and $\bm{(C_{2})}$ hold. This is analogous to imaginarity or coherence of quantum states. Since Choi states can embed quantum states, we also think that imaginarity measures of quantum states can also be used for quantum channels and present three imaginarity measures for quantum channels, they are $C_{r}$, $C_{t}$ and $C_{\alpha}$ via on the robustness, the trace norm and the sandwiched Rényi relative entropy, respectively.
Acknowledgement
We thank Jianwei Xu for helpful discussions and comments. This work was supported by National Natural Science Foundation of China (Grants No. 12271474).