### v2

# Imaginarity of quantum channels: Refinement and Alternative

Abstract
In this paper, we introduce the framework for quantifying the imaginarity of quantum channels. Besides, an alternative framework is given together to simplify the process of verifying the condition. We present three imaginarity measures of quantum channels based on the robustness, the trace norm, and the entropy, respectively. Some properties are also given.
keywords:
quantum resource theories , quantum channels , imaginarity measures , alternative framework
[…]
organization=School of Mathematics, addressline=Harbin Institute of Technology, city=Harbin, postcode=150001, country=China
1 Introduction
Imaginary numbers offer a broad perspective and provide additional methodologies for modern scientific research. Quantum resource theory facilitates the establishment of new connections across various research directions. In recent years, the concept of imaginarity resource theory [1], which has emerged alongside the development of quantum resource theory, has garnered increasing attention. Several applications can illustrate its significance. For instance, imaginarity plays a critical role in state discrimination [2]. The imaginarity operation proves to be particularly advantageous for distinguishing quantum channels without the need for an auxiliary system [3]. Numerous imaginarity measures have been identified for quantum states, including those based on fidelity [2], the convex roof [4], geometric considerations [3], and some induced by entropies [5, 6]. Some of them establish relations due to their mathematical properties [6, 7]. Just as quantum states warrant attention, quantum channels also merit investigation. Recent studies have begun to explore the properties of quantum channels [8, 9, 10, 11, 12]. The resource theories of quantum channels were first established in [13], with several subsequent theories proposed, such as coherence in quantum channels [14], entanglement in quantum channels [15], and more [16, 17]. Building upon these frameworks, additional measures for quantum channels have been introduced [18, 19, 20, 21].
The concept of the imaginarity of quantum channels was first introduced in [24], and the definitions of free channels and free superchannels have been revisited recently [25]. This indicates that the concept is of significant interest and worth considering. The structure of this paper is as follows: Similar to the coherence of quantum channels, we refine the concept of imaginarity for quantum channels and present an alternative framework that facilitates the verification of whether a given quantifier can serve as a suitable measure. We discuss the properties of imaginarity measures of quantum channels. Additionally, we propose three measures of the imaginarity of quantum channels based on the robustness, the trace norm, and the entropy, respectively.
2 Framework for quantifying the imaginarity of quantum channels
Let $H_{A}$ and $H_{B}$ be two Hilbert space with dimensions $|A|$ and $|B|$, and orthonormal basis $\left\{|j\rangle\right\}_{j=0}^{|A|-1}=\left\{|k\rangle\right\}_{k=0}^{|A|-1}$ and $\left\{|\alpha\rangle\right\}_{\alpha=0}^{|B|-1}=\left\{|\beta\rangle\right\}_{\beta=0}^{|B|-1}$. We assume $|a\rangle\langle a|=E_{a,a}$ without loss of generality, that is, it happens to be the a-th diagonal. Denote $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$ be the set of all density operators on $H_{A}$ and $H_{B}$, and $\mathcal{C}_{AB}$ be the set of all channels from $\mathcal{D}_{A}$ to $\mathcal{D}_{B}$. A quantum channel $\phi\in\mathcal{C}_{AB}$ can be represented by Choi matrix,
$J_{\phi}$ $=$ $\sum_{j,k}|j\rangle\langle k|\otimes\phi(|j\rangle\langle k|)$
$=$ $\sum_{j,k,\alpha,\beta}\phi_{j,k,\alpha,\beta}|j\rangle\langle k|\otimes|\alpha\rangle\langle\beta|$
with $\phi_{j,k,\alpha,\beta}=\langle\alpha|\phi(|j\rangle\langle k|)|\beta\rangle$.
The Choi state of $\phi$ is $\frac{J_{\phi}}{|A|}\triangleq\chi_{\phi}$. We call channel $\phi$ a real channel if $\chi_{\phi}$ is a real matrix, and denote $\mathcal{RC}_{AB}$ the set of all the real channels. The free channel is defined as the real channel. This is exactly the free operation in imaginarity resource theory, thus it is appropriate to define the free channel by Choi state.
Let $\mathcal{SC}_{ABA^{\prime}B^{\prime}}$ be the sets of all superchannels $\Theta$ from $\mathcal{C}_{AB}$ to $\mathcal{C}_{A^{\prime}B^{\prime}}$. The superchannel $\Theta$ can be represented by Choi matrix [14],
$J_{\Theta}$ $=$ $\sum_{j,k,\alpha,\beta}|j\alpha\rangle\langle k\beta|\otimes\Theta(|j\alpha\rangle\langle k\beta|)$
$=$ $\sum_{j,k,\alpha,\beta}\Theta_{j,k,\alpha,\beta,j^{\prime},k^{\prime},\alpha^{\prime},\beta^{\prime}}|j\alpha j^{\prime}\alpha\rangle\langle k\beta k^{\prime}\beta^{\prime}|,$
where $\Theta_{j,k,\alpha,\beta,j^{\prime},k^{\prime},\alpha^{\prime},\beta^{\prime}}=\langle j^{\prime}\alpha^{\prime}|\Theta(|j\alpha\rangle\langle k\beta|)|k^{\prime}\beta^{\prime}\rangle$ with $|j\alpha\rangle$ means the tensor of the basis from systems $A$ and $B$.
Similar to the case of channel, superchannel $\Theta$ also has the expression of Kraus operators $\left\{M_{m}\right\}_{m}$ with $\chi_{\Theta(\phi)}=\sum_{m}M_{m}\chi_{\phi}M_{m}^{\dagger}\triangleq\tilde{\Theta}(\chi_{\phi})$, where
$M_{m}=\sum_{j,j^{\prime},\alpha,\alpha^{\prime}}M_{m,j^{\prime},j,\alpha^{\prime},\alpha}|j^{\prime}\alpha\rangle\langle j\alpha|$
with $\Theta_{j,k,\alpha,\beta,j^{\prime},k^{\prime},\alpha^{\prime},\beta^{\prime}}=\sum_{m}M_{m,j^{\prime},j,\alpha^{\prime},\alpha}M_{m,k^{\prime},k,\beta^{\prime},\beta}$.
Here, $\tilde{\Theta}$ means the channel corresponding to the superchannel [14, 26]. We will not distinguish the signs, $\Theta$ or $\tilde{\Theta}$, since it can be judged by the context.
We call a superchannel $\Theta:\mathcal{C}_{AB}\longrightarrow\mathcal{C}_{A^{\prime}B^{\prime}}$ a real superchannels if $\Theta(\phi)\in\mathcal{RC}_{A^{\prime}B^{\prime}}$ for any channels $\phi\in\mathcal{RC}_{AB}$. We denote the set of all the real superchannels by $\mathcal{RSC}_{ABA^{\prime}B^{\prime}}$. The free superchannel is defined as the real superchannel.
In [24], the author presented a framework for quantifying the imaginarity of quantum channels. They introduced the imaginarity measure $C(\phi)\triangleq C(\chi_{\phi})$ of quantum channels should satisfy the following conditions:
$\bm{(C_{1})}$. Faithfulness: $C(\phi)\geqslant 0$, and $C(\phi)=0$ if and only if $\phi\in\mathcal{RC}_{AB}$;
$\bm{(C_{2})}$. Monotonicity: $C(\phi)\geqslant C(\Theta(\phi))$ for any real superchannel $\Theta\in\mathcal{RSC}_{ABA^{\prime}B^{\prime}}$.
Actually, strong monotonicity and convexity often have a great use and a proper quantum resource theory should include them. So we add these to the conditions above, and then the imaginarity measure $C$ of quantum channels becomes more familiar and complete.
$\bm{(C_{3})}$. Strong monotonicity: Let $\Theta\in\mathcal{RSC}_{ABA^{\prime}B^{\prime}}$ with $\Theta(\cdot)=\sum_{m}M_{m}\cdot M_{m}^{\dagger}$, then $C(\phi)\geqslant\sum_{m}p_{m}C(\phi_{m})$ where $p_{m}=\text{Tr}(M_{m}\chi_{\phi}M_{m}^{\dagger})$ and $\chi_{\phi_{m}}=M_{m}\chi_{\phi}M_{m}^{\dagger}/p_{m}$
$\bm{(C_{4})}$. Convexity: For any $\left\{\phi_{m}\right\}\subset\mathcal{C}_{AB}$ and any probability distribution $\left\{p_{m}\right\}$, $C(\sum_{m}p_{m}\phi_{m})\leqslant\sum_{m}p_{m}C(\phi_{m})$.
We consider that a proper imaginarity measure of quantum channels should satisfy the condition $\bm{(C_{1})}$$\bm{(C_{4})}$. In this way, we can obtain some properties of the imaginarity measure $C$.
Property 1.
Monotonicity under composition with any real channels:
$(a)$. Left composition: $C(\psi\circ\phi)\leqslant C(\phi)$ for any $\psi\in\mathcal{RC}_{BD}$ and any $\phi\in\mathcal{C}_{AB}$.
$(b)$. Right composition: $C(\phi\circ\psi)\leqslant C(\phi)$ for any $\psi\in\mathcal{RC}_{AB}$ and any $\phi\in\mathcal{C}_{BD}$.
$(C)$. Right composition: $C(\phi\circ\psi)=C(\phi)$ for any invertible $\psi\in\mathcal{RC}_{AB}$ and any $\phi\in\mathcal{C}_{BD}$.
Proof.
We only prove $(a)$, because $(b)$ is analogous. $(c)$ can be obtained from $(a)$ and $(b)$.
For $\psi\in\mathcal{RC}_{AB}$ and $\phi\in\mathcal{C}_{BD}$, we consider a superchannel $\Theta_{\psi}$ with $\Theta_{\psi}=\psi\circ\phi$. When superchannel $\Theta_{\psi}$ acts on any real channels, it is impossible to generate the imaginarity resource since the channel $\psi$ is real. It can be seen that the superchannel $\Theta_{\psi}$ is a real superchannel from the definition. The result holds due to the condition $\bm{(C_{2})}$. ∎
Property 2.
The combination of imaginarity measures is still an imaginarity measure, that is, for any $\left\{C_{j}\right\}$ and any probability distribution $\left\{q_{j}\right\}$ with $q_{j}\geqslant 0$, $C\triangleq\sum_{j}q_{j}C_{j}$ is an imaginarity measure.
Proof.
It is obvious that $C(\phi)\geqslant 0$. Due to $q_{j}\geqslant 0$, $C(\phi)=0$ is equivalent to every $C_{j}(\phi)=0$, and it indicate $\phi\in\mathcal{RC}$. For any real superchannel $\Theta$, we have $C(\phi)=\sum_{j}q_{j}C_{j}(\phi)\geqslant\sum_{j}q_{j}C_{j}(\Theta(\phi))=C(\Theta(\phi))$.
For $\Theta\in\mathcal{RSC}$ with $\Theta(\cdot)=\sum_{m}M_{m}\cdot M_{m}^{\dagger}$, the strong monotonicity holds:
$C(\phi)$
$=$ $\sum_{j}q_{j}C_{j}(\phi)$
$\geqslant$ $\sum_{j}q_{j}\sum_{m}p_{m}C_{j}(\phi_{m})$
$=$ $\sum_{m}p_{m}\sum_{j}q_{j}C_{j}(\phi_{m})$
$=$ $\sum_{m}p_{m}C(\phi_{m}).$
For any $\left\{\phi_{m}\right\}\subset\mathcal{C}_{AB}$ and any probability distribution $\left\{p_{m}\right\}$, the convexity holds:
$C(\sum_{m}p_{m}\phi_{m})$
$=$ $\sum_{j}q_{j}C_{j}(\sum_{m}p_{m}\phi_{m})$
$\leqslant$ $\sum_{j}q_{j}\sum_{m}p_{m}C_{j}(\phi_{m})$
$=$ $\sum_{m}p_{m}\sum_{j}q_{j}C_{j}(\phi_{m})$
$=$ $\sum_{m}p_{m}C(\phi_{m}).$
Therefore, $C$ is an imaginarity measure. ∎
Property 3.
If $C$ is an imaginarity measure for channel $\phi$, then the quantifier $\hat{C}$ as
$\hat{C}(\phi)\triangleq\min_{\left\{q_{j},\phi_{j}\right\}}\sum_{j}q_{j}C(\phi_{j}),$
is a proper measure for $\phi$, where $\min_{\left\{q_{j},\phi_{j}\right\}}$ represents the min of all possible channels decompositions as $\phi=\sum_{j}q_{j}\phi_{j}$ with probability distribution $q_{j}\geqslant 0$.
Proof.
It can be seen that $\hat{C}(\phi)\geqslant 0$. When $\hat{C}(\phi)=0$, we can take the best decomposition of $\phi$, expressed as $\phi=\sum_{k}q_{k}\phi_{k}$, so that $\hat{C}(\phi)=\sum_{k}q_{k}C(\phi_{k})$. We can get $C(\phi_{k})=0$ since $q\geqslant 0$, that is, $\phi_{k}\in\mathcal{RC}$, thus $\phi\in\mathcal{RC}$, the faithfulness holds.
For any probability distribution $\left\{q_{j}\right\}$, if we suppose every $\phi_{j}=\phi$, we can have $\sum_{j}q_{j}C(\phi_{j})=C(\phi)$. Thus, it follows that $\hat{C}(\phi)=\min_{\left\{q_{j},\phi_{j}\right\}}\sum_{j}q_{j}C(\phi_{j})\leqslant C(\phi)$. For any quantum channel $\phi$, we take the best decomposition of $\phi$, expressed as $\phi=\sum_{k}q_{k}\phi_{k}$, so that $\hat{C}(\phi)=\sum_{k}q_{k}C(\phi_{k})$. For $\Theta\in\mathcal{RSC}$ with $\Theta(\cdot)=\sum_{m}M_{m}\cdot M_{m}^{\dagger}$, we denote $a_{k,m}=\text{Tr}(M_{m}\chi_{\phi_{k}}M_{m}^{\dagger})$, $p_{m}=\text{Tr}(M_{m}\chi_{\phi}M_{m}^{\dagger})$, then
$\hat{C}(\phi)$
$=$ $\sum_{k}q_{k}C(\phi_{k})$
$\geqslant$ $\sum_{k}q_{k}\sum_{m}a_{k,m}C(\frac{M_{m}\chi_{\phi_{k}}M_{m}^{\dagger}}{a_{k,m}})$
$=$ $\sum_{m}p_{m}\sum_{k}\frac{q_{k}a_{k,m}}{p_{m}}C(\frac{M_{m}\chi_{\phi_{k}}M_{m}^{\dagger}}{a_{k,m}})$
$\geqslant$ $\sum_{m}p_{m}C(\sum_{k}\frac{q_{k}a_{k,m}}{p_{m}}\frac{M_{m}\chi_{\phi_{k}}M_{m}^{\dagger}}{a_{k,m}})$
$=$ $\sum_{m}p_{m}C(\frac{\sum_{k}q_{k}M_{m}\chi_{\phi_{k}}M_{m}^{\dagger}}{p_{m}})$
$=$ $\sum_{m}p_{m}C(\frac{M_{m}\chi_{\phi}M_{m}^{\dagger}}{p_{m}})$
$\geqslant$ $\sum_{m}p_{m}\hat{C}(\frac{M_{m}\chi_{\phi}M_{m}^{\dagger}}{p_{m}}).$
The first inequality follows from the strong monotonicity of $C$. The second inequality arises from the convexity of $C$ and the fact that $\sum_{k}\frac{q_{k}a_{k,m}}{p_{m}}=\frac{\sum_{k}q_{k}\text{Tr}(M_{m}\chi_{\phi_{k}}M_{m}^{\dagger})}{\text{Tr}(M_{m}\chi_{\phi}M_{m}^{\dagger})}=1$. The third inequality results from $\hat{C}(\phi)\leqslant C(\phi)$.
To verify the convexity of $\hat{C}$, for simplicity and without loss of generality, we consider only two quantum channels $\phi_{1}$ and $\phi_{2}$. We define $\psi=a\phi_{1}+b\phi_{2}$ with $a+b=1$. We take the best decomposition of $\phi_{1}$ and $\phi_{2}$ respectively as $\phi_{1}=\sum_{m}q_{1,m}\phi_{1,m}$ and $\phi_{2}=\sum_{n}q_{2,n}\phi_{2,n}$ so that the following equations hold:
$\begin{cases}\hat{C}(\phi_{1})=\sum_{m}q_{1,m}C(\phi_{1,m}),\\ \hat{C}(\phi_{2})=\sum_{n}q_{2,n}C(\phi_{2,n}).\end{cases}$
Then, $a\hat{C}(\phi_{1})+b\hat{C}(\phi_{2})=\sum_{m}aq_{1,m}C(\phi_{1,m})+\sum_{n}bq_{2,n}C(\phi_{2,n})$.
For ease of writing, if $m\neq n$, we supplement the elements $\left\{0,\bm{0}\right\}$ to the smaller set so that both sets have the same number of elements. This gives $a\hat{C}(\phi_{1})+b\hat{C}(\phi_{2})=\sum_{m}aq_{1,m}C(\phi_{1,m})+bq_{2,m}C(\phi_{2,m})$. Next, we introduce the new symbols $\left\{q_{2m-1},\psi_{2m-1}\right\}$ and $\left\{q_{2m},\psi_{2m}\right\}$ to represent $\left\{aq_{1,m},\phi_{1,m}\right\}$ and $\left\{bq_{2,m},\phi_{2,m}\right\}$ respectively. Then we have
$a\hat{C}(\phi_{1})+b\hat{C}(\phi_{2})$
$=$ $\sum_{m}q_{2m-1}C(\psi_{2m-1})+q_{2m}C(\psi_{2m})$
$\geqslant$ $\min_{q_{j},\psi_{j}}\sum_{j}q_{j}C(\psi_{j})$
$=$ $\hat{C}(\psi)=\hat{C}(a\phi_{1}+b\phi_{2}).$
In this way, we confirm that $\hat{C}$ is a valid imaginarity measure. ∎
3 Alternative framework for the imaginarity of quantum channels
In Section 2, we introduce the framework. However, conditions $\bm{(C_{3})}$ and $\bm{(C_{4})}$ can sometimes be difficult to verify. In coherence resource theory, to simplify the process of verifying whether a quantifier qualifies as a coherence measure, an alternative framework for quantifying coherence has been proposed [22]. Similar alternative frameworks can also be established in many other resource theories, such as imaginarity resource theory [7], and coherence of quantum channels [23]. Inspired by [23], we Introduce the following condition:
$\bm{(C_{5})}$. Additivity: $C(\phi)=p_{1}C(\phi_{1})+p_{2}C(\phi_{2})$ for any $p_{1}$, $p_{2}$ satisfied $p_{1}+p_{2}=1$, any $\phi_{1}\in\mathcal{C}_{AB_{1}}$ and $\phi_{2}\in\mathcal{C}_{AB_{2}}$, where $\phi=p_{1}\phi_{1}\oplus p_{2}\phi_{2}$, $\phi\in\mathcal{C}_{AB}$ with $|B|=|B_{1}|+|B_{2}|$.
When conditions $\bm{(C_{1})}$ and $\bm{(C_{2})}$ hold, the conditions $\bm{(C_{3})}$ and $\bm{(C_{4})}$ are equivalent to the condition $\bm{(C_{5})}$. To prove this, we first state the following theorem:
Theorem 1.
$\bm{(C_{3})}$ and $\bm{(C_{4})}$ give rise to $\bm{(C_{5})}$ when condition $\bm{(C_{1})}$ and $\bm{(C_{2})}$ holds.
Proof.
STEP 1, obtain: $\bm{C(\phi)\geqslant p_{1}C(\tilde{\phi}_{1})+p_{2}C(\tilde{\phi}_{2})}$.
Let $Q_{1}=|0\rangle\langle 0|+\cdots+||B_{1}|-1\rangle\langle|B_{1}|-1|+||B|\rangle\langle|B||+\cdots+||B|+|B_{1}|-1\rangle\langle|B|+|B_{1}|-1|+|(|A|-1)|B|\rangle\langle(|A|-1)|B||+\cdots+|(|A|-1)|B|+|B_{1}|-1\rangle\langle(|A|-1)|B|+|B_{1}|-1|$,
$Q_{2}=||B_{1}|\rangle\langle|B_{1}||+\cdots+||B|-1\rangle\langle|B|-1|+||B|+|B_{1}|\rangle\langle|B|+|B_{1}||+\cdots+||B|+|B|-1\rangle\langle|B|+|B|-1|+|(|A|-1)|B|+|B_{1}|\rangle\langle(|A|-1)|B|+|B_{1}||+\cdots+|(|A|-1)|B|+|B|-1\rangle\langle(|A|-1)|B|+|B|-1|$.
We can give an intuitive expression.
$Q_{1}=\overbrace{\begin{pmatrix}I_{|B_{1}|}&&&&\\ &O_{|B_{2}|}&&&\\ &&\ddots&&\\ &&&I_{|B_{1}|}&\\ &&&&O_{|B_{2}|}\end{pmatrix}}^{|A||B|}$
$=\bigoplus\limits_{j=0}^{|A|-1}\begin{pmatrix}I_{|B_{1}|}&\\ &O_{|B_{2}|}\end{pmatrix},$
$Q_{2}=\underbrace{\begin{pmatrix}O_{|B_{1}|}&&&&\\ &I_{|B_{2}|}&&&\\ &&\ddots&&\\ &&&O_{|B_{1}|}&\\ &&&&I_{|B_{2}|}\end{pmatrix}}_{|A||B|}$
$=\bigoplus\limits_{j=0}^{|A|-1}\begin{pmatrix}O_{|B_{1}|}&\\ &I_{|B_{2}|}\end{pmatrix},$
where $I_{a}$ and $O_{b}$ denote identity matrix and zero matrix with dimensions $a$ and $b$ respectively.
Given that $Q_{1}^{\dagger}Q_{1}+Q_{2}^{\dagger}Q_{2}=I_{|A||B|}$ and $Q_{1}Q_{2}=O_{|A||B|}$, let $\Theta(\cdot)=Q_{1}\cdot Q_{1}^{\dagger}+Q_{2}\cdot Q_{2}^{\dagger}$. It follows that $\Theta\in\mathcal{RSC}_{ABAB}$.
Under the requirement of $\bm{(C_{5})}$, the Choi state of $\phi$ is
$\chi_{\phi}$
$=$ $\sum_{j,k}|j\rangle\langle k|\otimes\phi(|j\rangle\langle k|)/|A|$
$=$ $\sum_{j,k}|j\rangle\langle k|\otimes\big{[}p_{1}\phi_{1}(|j\rangle\langle k|)\oplus p_{2}\phi_{2}(|j\rangle\langle k|\big{]}/|A|.$
Thus,
$Q_{1}\chi_{\phi}Q_{1}^{\dagger}=\sum_{j,k}|j\rangle\langle k|\otimes\big{[}p_{1}\phi_{1}(|j\rangle\langle k|)\oplus\bm{0}_{2}(|j\rangle\langle k|\big{]}/|A|.$
It is clear that $\text{Tr}(Q_{1}\chi_{\phi}Q_{1}^{\dagger})$ is exactly $p_{1}$. Let $\chi_{\tilde{\phi}_{1}}=Q_{1}\chi_{\phi}Q_{1}^{\dagger}/p_{1}=\sum_{j,k}|j\rangle\langle k|\otimes\big{[}\phi_{1}(|j\rangle\langle k|)\oplus\bm{0}_{2}(|j\rangle\langle k|\big{]}/|A|$. Similarly we can obtain $\chi_{\tilde{\phi}_{2}}$. From the condition $\bm{(C_{3})}$, we have:
$C(\phi)=C(\chi_{\phi})$$\geqslant$ $p_{1}C(\chi_{\tilde{\phi}_{1}})+p_{2}C(\chi_{\tilde{\phi}_{2}})$ (1)
$=$ $p_{1}C(\tilde{\phi}_{1})+p_{2}C(\tilde{\phi}_{2}).$
STEP 2, obtain: $\bm{C(\phi)\leqslant p_{1}C(\tilde{\phi}_{1})+p_{2}C(\tilde{\phi}_{2})}$.
For any $\rho\in\mathcal{D}_{A}$, $\rho=\sum_{j,k}\rho_{j,k}|j\rangle\langle k|$, any $\phi\in\mathcal{C}_{AB}$, $\phi(\rho)$ can be represented by $J_{\phi}$, in order to do this, $\rho$ and $J_{\phi}$ need to be tensor index realigned:
$\begin{cases}J^{\prime}_{\phi}=\sum_{j,k}\langle jk|\otimes\phi(|j\rangle\langle k|),\\ \rho^{\prime}=\sum_{j,k}\rho_{j,k}|jk\rangle,\end{cases}$
then $\phi(\rho)=J^{\prime}_{\phi}\rho^{\prime}=\sum_{j,k}\rho_{j,k}\phi(|j\rangle\langle k|)$. Under the requirement of $(C_{5})$, for any $\rho$ we have
$\phi(\rho)$
$=$ $\sum_{j,k}\rho_{j,k}\phi(|j\rangle\langle k|)$
$=$ $\sum_{j,k}\rho_{j,k}\big{[}p_{1}\phi_{1}(|j\rangle\langle k|)\oplus p_{2}\phi_{2}(|j\rangle\langle k|)\big{]}$
$=$ $p_{1}\sum_{j,k}\rho_{j,k}\big{[}\phi_{1}(|j\rangle\langle k|)\oplus\bm{0}_{2}(|j\rangle\langle k|)\big{]}$
$+p_{2}\sum_{j,k}\rho_{j,k}\big{[}\bm{0}_{1}(|j\rangle\langle k|)\oplus\phi_{2}(|j\rangle\langle k|)\big{]}$
$=$ $p_{1}\tilde{\phi}_{1}(\rho)+p_{2}\tilde{\phi}_{2}(\rho),$
that is $\phi=p_{1}\tilde{\phi}_{1}+p_{2}\tilde{\phi}_{2}$. From $\bm{(C_{4})}$, it follows that
$C(\phi)\leqslant p_{1}C(\tilde{\phi}_{1})+p_{2}C(\tilde{\phi}_{2}).$ (2)
STEP 3, obtain $\bm{(C_{5})}$ since $\bm{C(\phi_{j})=C(\tilde{\phi}_{j})}$, $\bm{j=1,2}$.
From inequalities (1) and (2) above, we obtain $C(\phi)=p_{1}C(\tilde{\phi}_{1})+p_{2}C(\tilde{\phi}_{2})$. Next, we need to prove that $C(\phi_{1})=C(\tilde{\phi}_{1})$, which requires us to show that $C(\chi_{\phi_{1}})=C(\chi_{\tilde{\phi}_{1}})$.
First, we specify the range of indices $i$ and $j$, let $i=0,1,\cdots,|B|-1$, $j=0,1,\cdots,|B_{1}|-1$. Let $K_{0}^{1}$ satisfies $\langle i|K_{0}^{1}|j\rangle=\delta_{i,j}$. An intuitive expression for $K_{0}^{1}$ is
$K_{0}^{1}=\left.\begin{pmatrix}I_{|B_{1}|}\\ \hat{O}\end{pmatrix}\right\}|B|,$
where $\hat{O}$ denotes the zero matrix with dimension $|B_{2}|\times|B_{1}|$.
Let $K_{n}^{2}$ satisfy $\langle j|K_{n}^{2}|i\rangle=\delta_{i,j+n|B_{1}|}$ for $n=0,1,\cdots,$ $\left\lceil|B_{2}|/|B_{1}|\right\rceil$, where $\left\lceil a\right\rceil$ means the ceil of $a$.
Let
$\mathcal{K}_{0}^{1}=\bigoplus_{j=0}^{|A|-1}K_{0}^{1}=\begin{pmatrix}K_{0}^{1}&&\\ &\ddots&\\ &&K_{0}^{1}\end{pmatrix},$
$\mathcal{K}_{n}^{2}=\bigoplus_{j=0}^{|A|-1}K_{n}^{2}=\begin{pmatrix}K_{n}^{2}&&\\ &\ddots&\\ &&K_{n}^{2}\end{pmatrix}.$
We have $(\mathcal{K}_{0}^{1})^{\dagger}\mathcal{K}_{0}^{1}=I_{|B_{1}|}$ and $\sum_{j=0}^{\left\lceil|B_{2}|/|B_{1}|\right\rceil}(\mathcal{K}_{n}^{2})^{\dagger}\mathcal{K}_{n}^{2}=I_{|B|}$. Thus $\Theta_{1}\in\mathcal{RSC}_{AB_{1}AB}$ with $\Theta_{1}(\cdot)=\mathcal{K}_{0}^{1}\cdot(\mathcal{K}_{0}^{1})^{\dagger}$, $\Theta_{2}\in\mathcal{RSC}_{ABAB_{1}}$ with $\Theta_{2}(\cdot)=\sum_{j=0}^{\left\lceil|B_{2}|/|B_{1}|\right\rceil}\mathcal{K}_{n}^{2}\cdot(\mathcal{K}_{n}^{2})^{\dagger}$. It can be seen that $\Theta_{1}(\chi_{\phi_{1}})=\chi_{\tilde{\phi}_{1}}$ and $\Theta_{2}(\chi_{\tilde{\phi}_{1}})=\chi_{\phi_{1}}$.
Combining with $\bm{(C_{2})}$, the above implies that the following inequality holds:
$C(\chi_{\phi_{1}})$
$=$ $C(\Theta_{2}(\chi_{\tilde{\phi}_{1}}))$
$\leqslant$ $C(\chi_{\tilde{\phi}_{1}})$
$=$ $C(\Theta_{1}(\chi_{\phi_{1}}))$
$\leqslant$ $C(\chi_{\phi_{1}}).$
As a result, $C(\chi_{\phi_{1}})=C(\chi_{\tilde{\phi}_{1}})$. It follows from a similar discussion that $C(\chi_{\phi_{2}})=C(\chi_{\tilde{\phi}_{2}})$.
Up to now, we prove $\bm{(C_{5})}$ holds: $C(\phi)=p_{1}C(\phi_{1})+p_{2}C(\phi_{2})$. ∎
Under the requirement of $\bm{(C_{5})}$, we have $\chi_{\phi_{1}}$ and $\chi_{\phi_{2}}$ below.
$\begin{cases}\chi_{\phi_{1}}=\sum_{j,k}|j\rangle\langle k|\otimes\phi_{1}(|j\rangle\langle k|)/|A|,\\ \chi_{\phi_{2}}=\sum_{j,k}|j\rangle\langle k|\otimes\phi_{2}(|j\rangle\langle k|)/|A|.\end{cases}$
We also define $\chi_{\tilde{\phi}_{1}}$ and $\chi_{\tilde{\phi}_{2}}$ below.
$\begin{cases}\chi_{\tilde{\phi}_{1}}=\sum_{j,k}|j\rangle\langle k|\otimes\big{[}\phi_{1}(|j\rangle\langle k|)\oplus\bm{0}_{2}(|j\rangle\langle k|\big{]}/|A|,\\ \chi_{\tilde{\phi}_{2}}=\sum_{j,k}|j\rangle\langle k|\otimes\big{[}\bm{0}_{1}(|j\rangle\langle k|)\oplus\phi_{2}(|j\rangle\langle k|\big{]}/|A|.\end{cases}$
Thus the following property holds.
Next, we prove the reverse result:
Theorem 2.
$\bm{(C_{5})}$ gives rise to $\bm{(C_{3})}$ and $\bm{(C_{4})}$ when condition $\bm{(C_{1})}$ and $\bm{(C_{2})}$ holds.
Proof.
prove $\bm{(C_{3})}$.
For $\Theta\in\mathcal{RSC}_{ABA^{\prime}B^{\prime}}$ with $\Theta(\cdot)=\sum_{m}M_{m}\cdot M_{m}^{\dagger}$, we introduce an auxiliary system $S$ with dimension $N$ and orthonormal basis $\left\{|n\rangle\right\}_{n=0}^{N-1}$. The system $AB$, formed by Choi state and the auxiliary system $S$, constitutes a combined system $ABS$, whose basis is $\left\{|j\alpha\rangle\otimes|n\rangle\right\}$. The entire system is initially in the state $\chi_{\phi}\otimes|0\rangle\langle 0|\triangleq\chi_{\phi}^{ABS}$.
Let $U_{n}=\sum_{k=0}^{N-1}|(k+n)\text{ mod }N\rangle\langle k|$, then $\Theta^{ABS}(\cdot)\in\mathcal{RSC}_{ABSA^{\prime}B^{\prime}S}$ with $\Theta^{ABS}(\cdot)=\sum_{n}(M_{n}\otimes U_{n})\cdot(M_{n}\otimes U_{n})^{\dagger}$. Thus
$\Theta^{ABS}(\chi_{\phi}^{ABS})$
$=$ $\sum_{n}(M_{n}\otimes U_{n})(\chi_{\phi}\otimes|0\rangle\langle 0|)(M_{n}\otimes U_{n})^{\dagger}$
$=$ $\sum_{n}M_{n}\chi_{\phi}M_{n}^{\dagger}\otimes|n\rangle\langle n|$
$=$ $\sum_{n}p_{n}\chi_{\phi_{n}}\otimes|n\rangle\langle n|,$
where $p_{n}=\text{Tr}(M_{n}\chi_{\phi}M_{n}^{\dagger})$ and $\chi_{\phi_{n}}=M_{n}\chi_{\phi}M_{n}^{\dagger}/p_{n}$.
It can be seen that $\sum_{n}p_{n}=1$, $\phi_{n}\in\mathcal{C}_{AB_{n}}$ with $|B|=\sum_{n}|B_{n}|$. From $\bm{(C_{5})}$, we have
$C(\chi_{\phi}^{ABS})=C(\chi_{\phi}\otimes|0\rangle\langle 0|)=C(\chi_{\phi}),$
and
$C(\Theta^{ABS}(\chi_{\phi}^{ABS}))$
$=$ $C(\sum_{n}p_{n}\chi_{\phi_{n}}\otimes|n\rangle\langle n|)$
$=$ $C(\bigoplus_{n}p_{n}\chi_{\phi_{n}})$
$=$ $\sum_{n}p_{n}C(\chi_{\phi_{n}}).$
From $\bm{(C_{2})}$ and above, we can obtain
$C(\chi_{\phi})\geqslant\sum_{n}p_{n}C(\chi_{\phi_{n}}),$
that is, condition $\bm{(C_{3})}$ holds.
prove $\bm{(C_{4})}$.
For any $\left\{\phi_{m}\right\}\subset\mathcal{C}_{AB}$ and any probability distribution $\left\{p_{m}\right\}$, we construct a auxiliary system $S$ with orthonormal basis $\left\{|m\rangle\right\}_{m}$ and define $\chi_{\phi}^{ABS}=\sum_{m}p_{m}\chi_{\phi_{m}}\otimes|m\rangle\langle m|$. Notably, we have:
$\sum_{m}(I_{|A||B|}\otimes|0\rangle\langle m|)^{\dagger}(I_{|A||B|}\otimes|0\rangle\langle m|)=I_{|A||B||S|}$
and $\Theta^{ABS}\in\mathcal{RSC}_{ABSA^{\prime}B^{\prime}S}$ with $\Theta^{ABS}(\cdot)=\sum_{m}(I_{|A||B|}\otimes|0\rangle\langle m|)\cdot(I_{|A||B|}\otimes|0\rangle\langle m|)^{\dagger}$. Then,
$\Theta^{ABS}(\chi_{\phi}^{ABS})$
$=$ $\sum_{m}(I_{|A||B|}\otimes|0\rangle\langle m|)\chi_{\phi}^{ABS}(I_{|A||B|}\otimes|0\rangle\langle m|)^{\dagger}$
$=$ $\sum_{m}p_{m}\chi_{\phi_{m}}\otimes|0\rangle\langle 0|.$
From $\bm{(C_{2})}$ and $\bm{(C_{5})}$, the following inequality holds:
$\sum_{m}p_{m}C(\chi_{\phi_{m}})$
$=$ $C(\chi_{\phi}^{ABS})$
$\geqslant$ $C(\Theta^{ABS}(\chi_{\phi}^{ABS}))$
$=$ $C(\sum_{m}p_{m}\chi_{\phi_{m}}),$
that is, condition $\bm{(C_{4})}$ holds. ∎
Therefore, from the above two theorems, we can give the Alternative framework for the imaginarity of quantum channels:
Theorem 3.
$\bm{(C_{3})}+\bm{(C_{4})}$ is equivalent to $\bm{(C_{5})}$ when condition $\bm{(C_{1})}$ and $\bm{(C_{2})}$ holds.
When verifying whether a quantifier is an imaginarity measure, we can choose to verify either the conditions $\bm{(C_{1})}$$\bm{(C_{4})}$ or the conditions $\bm{(C_{1})}+\bm{(C_{2})}+\bm{(C_{5})}$.
Property 4.
$C$ is a proper imaginarity measure for quantum channels, the equation holds:
$C(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})=C(p_{1}\chi_{\tilde{\phi}_{1}}+p_{2}\chi_{\tilde{\phi}_{2}})=C(\chi_{\phi}).$ (3)
Proof.
We employ the same idea as in Theorem 1, specifically using $\bm{(C_{2})}$ to establish equivalence. Without loss of generality, we assume $|A|=3$, $p_{1}\phi_{1}(|j\rangle\langle k|)\triangleq X_{jk}$ and $p_{2}\phi_{2}(|j\rangle\langle k|)\triangleq Y_{jk}$. Then,
$p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}}=\begin{pmatrix}X_{11}&X_{12}&X_{13}&&&\\ X_{21}&X_{22}&X_{23}&&&\\ X_{31}&X_{32}&a_{33}&&&\\ &&&Y_{11}&Y_{12}&Y_{13}\\ &&&Y_{21}&Y_{22}&Y_{23}\\ &&&Y_{31}&Y_{32}&Y_{33}\end{pmatrix},$
and
$p_{1}\chi_{\tilde{\phi}_{1}}+p_{2}\chi_{\tilde{\phi}_{2}}=\begin{pmatrix}X_{11}&&X_{12}&&X_{13}&\\ &Y_{11}&&Y_{12}&&Y_{13}\\ X_{21}&&X_{22}&&X_{23}&\\ &Y_{21}&&Y_{22}&&Y_{23}\\ X_{31}&&X_{32}&&X_{33}&\\ &Y_{31}&&Y_{32}&&Y_{33}\end{pmatrix}.$
In this way, we need to find some superchannels between $p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}}$ and $\chi_{\phi}$. We can set
$K_{1}=\begin{pmatrix}I_{|B_{1}|}&&&&&\\ &&&I_{|B_{1}|}&&\\ &&&&I_{|B_{1}|}&\\ &I_{|B_{2}|}&&&&\\ &&I_{|B_{2}|}&&&\\ &&&&&I_{|B_{2}|}\end{pmatrix},$
$K_{2}=\begin{pmatrix}I_{|B_{1}|}&&&&&\\ &I_{|B_{2}|}&&&&\\ &&I_{|B_{1}|}&&&\\ &&&I_{|B_{2}|}&&\\ &&&&I_{|B_{1}|}&\\ &&&&&I_{|B_{2}|}\end{pmatrix}.$
It can be known that $K^{\dagger}_{j}K_{j}=I$, $j=1,2$, and the following transformations,
$p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}}\xrightarrow{K_{1}(\cdot)K_{1}^{\dagger}}N\xrightarrow{K_{2}(\cdot)K_{2}^{\dagger}}p_{1}\chi_{\tilde{\phi}_{1}}+p_{2}\chi_{\tilde{\phi}_{2}},$
where $N$ is
$\begin{pmatrix}X_{11}&&&X_{12}&X_{13}&\\ &Y_{11}&Y_{12}&&&Y_{13}\\ &Y_{21}&Y_{22}&&&Y_{23}\\ X_{21}&&&X_{22}&X_{23}&\\ X_{31}&&&X_{32}&X_{33}&\\ &Y_{31}&Y_{32}&&&Y_{33}\end{pmatrix}.$
We define $\Theta_{j}(\cdot)=K_{j}\cdot K_{j}^{\dagger}$, $j=1,2$. It is clear that $\Theta_{j}\in\mathcal{RSC}$. Then the inequality holds:
$C(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})\geqslant C(p_{1}\chi_{\tilde{\phi}_{1}}+p_{2}\chi_{\tilde{\phi}_{2}}).$
Since $K_{1}$ and $K_{2}$ are invertible, the inverse inequality also holds. Therefore, we have:
$C(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})=C(p_{1}\chi_{\tilde{\phi}_{1}}+p_{2}\chi_{\tilde{\phi}_{2}}).$
When the dimension $|A|$ takes other values, we can similarly construct many invertible real superchannels such that
$\Theta_{1}(\cdots\Theta_{k}(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})\cdots)=p_{1}\chi_{\tilde{\phi}_{1}}+p_{2}\chi_{\tilde{\phi}_{2}}.$
In this way, we can conclude the equation (3) holds. ∎
4 Imaginarity measures of quantum channels
In this section, we introduce three imaginarity measures of quantum channels, that is, measures based on the robustness, the trace norm, and the sandwiched Rényi relative entropy, respectively.
4.1 Measure based on the robustness
The Measure based on the robustness has appeared in many resource theories [1, 18]. Similarly, we define the imaginarity measure of quantum channels as follows.
$C_{r}(\phi)=\min_{\tilde{\phi}\in\mathcal{RC}_{AB}}\left\{s\geqslant 0:\frac{\chi_{\phi}+s\chi_{\tilde{\phi}}}{1+s}\triangleq\chi_{\psi},\psi\in\mathcal{RC}_{AB}\right\}$
We have the following result.
Theorem 4.
$C_{r}(\phi)$ is an imaginarity measure of quantum channels.
Before our proof, we need to introduce pseudomixture [27, 28, 29, 7] of $\chi_{\phi}$, that is,
$\chi_{\phi}=[1+C_{r}(\chi_{\phi})]\chi_{\phi_{1}}-C_{r}(\chi_{\phi})\chi_{\phi_{2}},$
where $\phi_{1}\in\mathcal{RC}_{AB}$, and $\phi_{2}$ is the optimal channel of the $\min$ in the definition of $C_{r}(\phi)$.
Proof.
We only verify $\bm{(C_{1})}$+$\bm{(C_{3})}$+$\bm{(C_{4})}$ since $\bm{(C_{2})}$ can be derived by $\bm{(C_{3})}$ and$\bm{(C_{4})}$.
$\bm{(C_{1})}$:
It can be observed that $C_{r}(\phi)\geqslant 0$. When $C_{r}(\phi)=0$, this implies $\frac{\chi_{\phi}+0}{1+0}=\chi_{\psi}$, which means $\phi=\psi\in\mathcal{RC}_{AB}$. The converse is also clearly true.
$\bm{(C_{3})}$:
Let $\Theta\in\mathcal{RSC}_{ABA^{\prime}B^{\prime}}$ with $\Theta(\cdot)=\sum_{m}M_{m}\cdot M_{m}^{\dagger}$. For any quantum channel $\phi$ with its pseudomixture, under the requirement of $\bm{(C_{3})}$, we have $p_{m}=\text{Tr}(M_{m}\chi_{\phi}M_{m}^{\dagger})$ and $\chi_{\phi_{m}}=M_{m}\chi_{\phi}M_{m}^{\dagger}/p_{m}$. Then, we obtain:
$M_{m}\chi_{\phi}M_{m}^{\dagger}$
$=$ $[1+C_{r}(\chi_{\phi})]M_{m}\chi_{\phi_{1}}M_{m}^{\dagger}-C_{r}(\chi_{\phi})M_{m}\chi_{\phi_{2}}M_{m}^{\dagger}.$
Let
$\begin{cases}\chi_{\phi_{1}}^{m}=\frac{1}{1+s_{m}}\frac{1}{p_{m}}[1+C_{r}(\chi_{\phi})]M_{m}\chi_{\phi_{1}}M_{m}^{\dagger},\\ \chi_{\phi_{2}}^{m}=\frac{1}{s_{m}}\frac{1}{p_{m}}C_{r}(\chi_{\phi})M_{m}\chi_{\phi_{2}}M_{m}^{\dagger},\end{cases}$
where $s_{m}=\frac{1}{p_{m}}C(\chi_{\phi})\text{Tr}(M_{m}\chi_{\phi_{1}}M_{m}^{\dagger})$.
We obtain $\chi_{\phi_{m}}=(1+s_{m})\chi_{\phi_{1}}^{m}-s_{m}\chi_{\phi_{2}}^{m}$. This means $\frac{\chi_{\phi_{m}}+s_{m}\chi_{\phi_{2}}^{m}}{1+s_{m}}=\chi_{\phi_{1}}^{m}$. Thus we get $C_{r}(\chi_{\phi_{m}})\leqslant s_{m}$ from definition, that is, $p_{m}C_{r}(\chi_{\phi_{m}})\leqslant C(\chi_{\phi})\text{Tr}(M_{m}\chi_{\phi_{1}}M_{m}^{\dagger})$.
Therefore, we have:
$\sum_{m}p_{m}C_{r}(\chi_{\phi_{m}})$
$\leqslant$ $\sum_{m}C(\chi_{\phi})\text{Tr}(M_{m}\chi_{\phi_{1}}M_{m}^{\dagger})$
$=$ $C(\chi_{\phi})\sum_{m}\text{Tr}(M_{m}^{\dagger}M_{m}\chi_{\phi_{1}})$
$=$ $C(\chi_{\phi})\text{Tr}(\sum_{m}M_{m}^{\dagger}M_{m}\chi_{\phi_{1}})$
$=$ $C(\chi_{\phi})$
This means the condition $\bm{(C_{3})}$ holds.
$\bm{(C_{4})}$:
For simplicity, we consider only two quantum channels $\phi_{1}$ and $\phi_{2}$, and define $\phi=p\phi_{1}+(1-p)\phi_{2}$ with any $p\in(0,1)$. This leads to $\chi_{\phi}=p\chi_{\phi_{1}}+(1-p)\chi_{\phi_{2}}$. Let $s_{p}=pC_{r}(\chi_{\phi_{1}})+(1-p)C_{r}(\chi_{\phi_{2}})$.
Consider the pseudomixtures of $\chi_{\phi_{1}}$ and $\chi_{\phi_{2}}$:
$\chi_{\phi_{1}}=[1+C_{r}(\chi_{\phi_{1}})]\chi_{\phi_{1,1}}-C_{r}(\chi_{\phi_{1}})\chi_{\phi_{1,2}},$
$\chi_{\phi_{2}}=[1+C_{r}(\chi_{\phi_{2}})]\chi_{\phi_{2,1}}-C_{r}(\chi_{\phi_{2}})\chi_{\phi_{2,2}},$
then let
$\begin{cases}\chi_{\psi}=\frac{p[1+C_{r}(\chi_{\phi_{1}})]\chi_{\phi_{1,1}}+(1-p)[1+C_{r}(\chi_{\phi_{2}})]\chi_{\phi_{2,1}}}{1+s_{p}},\\ \chi_{\tilde{\phi}}=\frac{pC_{r}(\chi_{\phi_{1}})\chi_{\phi_{1,2}}+(1-p)C_{r}(\chi_{\phi_{2}})\chi_{\phi_{2,2}}}{s_{p}}.\end{cases}$
We have $\chi_{\phi}=(1+s_{p})\chi_{\psi}-s_{p}\chi_{\tilde{\phi}}$, that is, $\chi_{\psi}=\frac{\chi_{\phi}+s_{p}\chi_{\tilde{\phi}}}{1+s_{p}}$, thus we obtain $s_{p}\geqslant C_{r}(\chi_{\phi})$.
Therefore, $pC_{r}(\chi_{\phi_{1}})+(1-p)C_{r}(\chi_{\phi_{2}})=s_{p}\geqslant C_{r}(\chi_{\phi})=C(p\chi_{\phi_{1}}+(1-p)\chi_{\phi_{2}})$. The condition $\bm{(C_{4})}$ holds. ∎
4.2 Measure based on the trace norm
The trace norm,
$C_{t}(\phi)=||\chi_{\phi}-\chi_{\phi}^{*}||_{\text{Tr}}=\text{Tr}|\chi_{\phi}-\chi_{\phi}^{*}|.$
can be used to construct a measure, where $\chi_{\phi}^{*}$ means the conjugate of $\chi_{\phi}$ and $|\chi_{\phi}-\chi_{\phi}^{*}|=\sqrt{(\chi_{\phi}-\chi_{\phi}^{*})^{\dagger}(\chi_{\phi}-\chi_{\phi}^{*})}$.
Theorem 5.
$C_{t}(\phi)$ is a proper imaginarity measure of quantum channels.
Proof.
The condition $\bm{(C_{1})}$ holds due to the definition of the trace norm.
Due to $||\chi_{\phi}-\chi_{\phi}^{*}||_{\text{Tr}}\geqslant||\Theta(\chi_{\phi})-\Theta(\chi_{\phi}^{*})||_{\text{Tr}}$ for any CPTP map [30], it certainly holds for real superchannels. Since $\Theta$ is a real superchannel, it can be seen that $[\Theta(\chi_{\phi})]^{*}=\Theta(\chi_{\phi}^{*})$ due to
$[\Theta(\chi_{\phi})]^{*}$
$=$ $\left[\sum_{m}M_{m}\chi_{\phi}M_{n}^{\dagger}\right]^{*}$
$=$ $\sum_{m}M_{m}\chi_{\phi}^{*}M_{n}^{\dagger}$
$=$ $\Theta(\chi_{\phi}^{*}).$
Thus $\bm{(C_{2})}$ is satisfied.
Under the requirement of $\bm{(C_{5})}$, from Property 4, we have:
$C_{t}(\chi_{\phi})$
$=$ $C_{t}(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})$
$=$ $||(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})-(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})^{*}||_{\text{Tr}}$
$=$ $||(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})-(p_{1}\chi_{\phi_{1}}^{*}\oplus p_{2}\chi_{\phi_{2}}^{*})||_{\text{Tr}}$
$=$ $||p_{1}(\chi_{\phi_{1}}-\chi_{\phi_{1}}^{*})\oplus p_{2}(\chi_{\phi_{2}}-\chi_{\phi_{2}}^{*})||_{\text{Tr}}$
$=$ $p_{1}||(\chi_{\phi_{1}}-\chi_{\phi_{1}}^{*})||_{\text{Tr}}+p_{2}||(\chi_{\phi_{2}}-\chi_{\phi_{2}}^{*})||_{\text{Tr}}$
$=$ $p_{1}C_{t}(\chi_{\phi_{1}})+p_{2}C_{t}(\chi_{\phi_{2}}).$
Therefore, $C_{r}(\phi)$ satisfies $\bm{(C_{1})}$+$\bm{(C_{2})}$+$\bm{(C_{5})}$. ∎
4.3 Measure based the sandwiched Rényi relative entropy
The sandwiched Rényi relative entropy [31] is defined for quantum states with the parameter $\alpha$ as follows:
$D_{\alpha}(\rho||\sigma)=\frac{1}{\alpha-1}\text{log }\text{Tr}(\sigma^{\frac{1-\alpha}{2\alpha}}\rho\sigma^{\frac{1-\alpha}{2\alpha}})^{\alpha},$
Now, we will introduce the imaginarity measure of quantum channels based on the sandwiched Rényi relative entropy, defined as:
$C_{\alpha}(\phi)=1-\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha},$
where $\frac{1}{2}\leqslant\alpha<1$. We have the following results:
Theorem 6.
$C_{\alpha}(\phi)$ is a proper imaginarity measure of quantum channels.
Proof.
It is shown [31] that $D_{\alpha}(\rho||\sigma)\geqslant 0$ when $\frac{1}{2}\leqslant\alpha<1$, and equation holds if and only if $\rho=\sigma$, which implies that $\text{Tr}(\sigma^{\frac{1-\alpha}{2\alpha}}\rho\sigma^{\frac{1-\alpha}{2\alpha}})^{\alpha}\leqslant 1$. Thus, considering Choi state as a quantum state, we find that $C_{\alpha}(\phi)\geqslant 0$. The equation holds if and only if $\chi_{\phi}=\chi_{\phi}^{*}$, satisfying the condition $\bm{(C_{1})}$.
From [31], we have:
$D_{\alpha}(\chi_{\phi}||\chi_{\phi}^{*})\geqslant D_{\alpha}\left[\Theta(\chi_{\phi})||\Theta(\chi_{\phi}^{*})\right]$
$\iff$ log $\text{log }\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}$
$\leqslant\text{log }\text{Tr}\left\{[\Theta(\chi_{\phi}^{*})]^{\frac{1-\alpha}{2\alpha}}\Theta(\chi_{\phi})[\Theta(\chi_{\phi}^{*})]^{\frac{1-\alpha}{2\alpha}}\right\}^{\alpha}$
$\iff$ $\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}$
$\leqslant\text{Tr}\left\{[\Theta(\chi_{\phi}^{*})]^{\frac{1-\alpha}{2\alpha}}\Theta(\chi_{\phi})[\Theta(\chi_{\phi}^{*})]^{\frac{1-\alpha}{2\alpha}}\right\}^{\alpha}$
$\iff$ $\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}$
$\leqslant\text{Tr}\left\{([\Theta(\chi_{\phi})]^{*})^{\frac{1-\alpha}{2\alpha}}\Theta(\chi_{\phi})([\Theta(\chi_{\phi})]^{*})^{\frac{1-\alpha}{2\alpha}}\right\}^{\alpha}$
$\iff$ $C_{\alpha}(\phi)\geqslant C_{\alpha}\left[\Theta(\phi)\right].$
Thus $\bm{(C_{2})}$ holds.
Under the requirement of $\bm{(C_{5})}$, we have:
$\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}$
$=$ $\text{Tr}\left\{\bigoplus_{j=1}^{2}p_{j}^{\frac{1-\alpha}{2\alpha}}p_{j}p_{j}^{\frac{1-\alpha}{2\alpha}}(\chi_{\phi_{j}}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi_{j}}(\chi_{\phi_{j}}^{*})^{\frac{1-\alpha}{2\alpha}}\right\}^{\alpha}$
$=$ $\sum_{j=1}^{2}p_{j}\text{Tr}\left[(\chi_{\phi_{j}}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi_{j}}(\chi_{\phi_{j}}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}$
In this way, we obtain:
$C_{\alpha}(\phi)$
$=$ $1-\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}$
$=$ $\sum_{j=1}^{2}p_{j}\left\{1-\text{Tr}\left[(\chi_{\phi_{j}}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi_{j}}(\chi_{\phi_{j}}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}\right\}$
$=$ $p_{1}C_{\alpha}(\phi_{1})+p_{2}C_{\alpha}(\phi_{2})$
The condition $\bm{(C_{5})}$ holds. ∎
5 Conclusion
We introduce the framework of imaginarity of quantum channels, and introduce an alternative framework. Besides, we give some properties of imaginarity measure of quantum channels. We also give three imaginarity measures for quantum channels. They are $C_{r}$, $C_{t}$ and $C_{\alpha}$ based on the robustness, the trace norm and the sandwiched Rényi relative entropy, respectively.
Acknowledgement
We thank Jianwei Xu for helpful discussions and comments. This work was supported by National Natural Science Foundation of China (Grants No. 12271474).